### v7

# Twisted Superfields

Abstract
A model is presented that could lead to an interesting extension of the Standard Model. Like a supersymmetric gauge theory, the model is holomorphic and invariant to local superspace gauge transformations. However, the model is not invariant to superspace translations, so it is not supersymmetric. It is proposed that this combination allows the model to have many of the attractive features of supersymmetric theories, while at the same time predicting fewer particles that have not yet been seen experimentally. For example, the “superpartners” of the gauge bosons in the model are quarks. The model is able to generate the symmetries and particles of the Standard Model, but with some significant differences that have observable consequences. These consequences provide possible explanations for a number of 3-7 sigma deviations from Standard Model calculations that have been found in recent experiments.
Introduction
Supersymmetric gauge theories have many attractive properties. For example, they are (i) holomorphic, (ii) invariant to local gauge transformations in superspace, and (iii) invariant to global translations in superspace. A difficulty with supersymmetric theories, however, is that they predict that for every particle that has been observed, there is another partner particle that has not yet been seen. As experiments probe higher and higher energies, the fact that no partner particle has been found becomes more problematic. Early on, supersymmetry practitioners asked whether some of the existing observed particles could actually be supersymmetric partners with each other. The HLS theorem [1] mostly rules out this possibility with some minor exceptions (like the Higgs boson being a slepton [2]). But even those exceptions are generally not accepted for other reasons.
The model presented in this paper incorporates the first two features of supersymmetry listed above, but not the third. Since the model is not invariant to superspace translations, it is not supersymmetric. Consequently, there is no a priori guarantee that quadratic divergences cancel. That being said, many quadratic divergences are cancelled for supersymmetric gauge theories primarily due to the fact that the theories are holomorphic and invariant to superspace gauge transformations. Also, local superspace gauge transformations turn scalar bosons into fermions, and fermions do not have quadratic divergences. Divergences may cancel in this theory for similar reasons. To that point, for nonsupersymmetric gauge theories similar to the one presented in this paper, it has been shown explicitly that quadratic divergences cancel to at least the two-loop order [3, 4].
The superspace gauge transformations of the model are built on the group U(3)$\times$×U(3). The field content of the model includes constructions built in N=1 superspace that are not N=1 superfields but are nonetheless called “twisted superfields” by way of analogy. Specifically, the model includes a real “twisted superfield”, an adjoint-representation chiral “twisted superfield”, and three flavors of fundamental and anti-fundamental chiral “twisted superfields”.
An advantage to this theory not being supersymmetric is that it can allow existing observed particles to be “superpartners” with each other (in the sense that superspace gauge transformations change them into each other). For example, in this model the partners of the gauge bosons (within the real “twisted superfield”) are quarks. Also, due to the fact that the theory is holomorphic and gauge invariant, it is argued that some of the nonperturbative phenomena of supersymmetric gauge theories may apply to the present theory.
[…]
Many of the ideas of this paper were originally published by the author in [6]. However, this paper has heavily revised the structure of the theory, the parameter values, and the mapping to experimental results. This paper replaces that original paper.
1 U(3) x U(3) Symmetries and Fields
The theory is constructed in N=1 superspace, extending four-dimensional spacetime by including four additional anticommuting coordinates $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$. For reviews of superspace, see [7, 8, 9, 10, 11, 12, 13]; the notational conventions of [7] are used throughout. The gauge group of the model is U(3)$\times$×U(3), and it is described using 6$\times$6 matrices with the gauge fields in the 3$\times$3 diagonal blocks. The model includes a construction called a real “twisted superfield” that has the following attributes: field components in the 3$\times$3 diagonal blocks have an even number of $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$ factors, while those in the 3$\times$3 off-diagonal blocks have an odd number of $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$ factors. An adjoint-representation chiral “twisted superfield” has this same structure, and fundamental chiral “twisted superfields” have structures consistent with those.
For normal superfields, translations in superspace transform a bosonic component of the superfield into a fermionic one, and vice versa. For the constructions used here to be superfields, every translation in superspace would have to be accompanied by a gauge rotation. But the HLS theorem has proven that this kind of rotation is not allowed for N=1 supersymmetric theories [1]. Therefore, the constructions used to develop the model are not N=1 superfields, but the term is used to provide an analogy with supersymmetric theories.
The real “twisted superfield” $V=V^{{\dagger}}$ is defined by:
$V=\left(\begin{array}[]{cc}{C_{1}+N_{1}\theta^{2}+\bar{\theta}^{2}N_{1}^{{\dagger}}-\bar{\theta}\bar{\sigma}^{\mu}A_{1\mu}\theta+{\tfrac{1}{2}}\bar{\theta}^{2}d_{1}\theta^{2}}&{\eta\theta+\bar{\theta}\tilde{\eta}^{{\dagger}}+i\bar{\theta}\tilde{\lambda}^{{\dagger}}\theta^{2}-i\bar{\theta}^{2}\lambda\theta}\\ {\tilde{\eta}\theta+\bar{\theta}\eta^{{\dagger}}+i\bar{\theta}\lambda^{{\dagger}}\theta^{2}-i\bar{\theta}^{2}\tilde{\lambda}\theta}&{C_{2}+N_{2}\theta^{2}+\bar{\theta}^{2}N_{2}^{{\dagger}}-\bar{\theta}\bar{\sigma}^{\mu}A_{2\mu}\theta+{\tfrac{1}{2}}\bar{\theta}^{2}d_{2}\theta^{2}}\end{array}\right),$ (1.1)
where each component field above is a U(3) matrix function of spacetime coordinates $x^{\mu}$. For example, $A_{1\mu}=A_{1\mu}^{A}\left(x\right)t^{A}$, where $t^{A}$ are 3$\times$3 U(3) matrices normalized by ${\rm tr}\left(t^{A}t^{B}\right)={\tfrac{1}{2}}\delta^{AB}$. Lower case letters are used to denote SU(3) adjoint indices $a,b\in\left\{1,2...8\right\}$. Upper case letters are used to denote U(3) adjoint indices $A,B\in\left\{0,1,2...8\right\}$ that include the Abelian matrix $t^{0}={\tfrac{1}{\sqrt{6}}}{\rm diag}\left(1,1,1\right)$. The $\theta_{\alpha}$ are 2-component anti-commuting Grassman coordinates, and $\bar{\theta}_{\dot{\alpha}}$ are their Hermitian conjugates. As a result of their $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$ factors, the fields in the diagonal blocks of $V$ are bosons, while the fields in the off-diagonal blocks are fermions.
As mentioned previously, a theory built using the above real twisted superfield is not supersymmetric, since the fermion fields in the superfield are in a different representation of the U(3)$\times$×U(3) group than the boson fields. Despite not being supersymmetric, the real twisted superfield is assumed to transform as follows under a local “twisted supergauge transformation”:
$e^{V}\to e^{i\Lambda^{{\dagger}}}e^{V}e^{-i\Lambda}.$ (1.2)
In the above expression,
$\Lambda=\left(\begin{array}[]{cc}{\alpha_{1}\left(y\right)+\theta^{2}n_{1}\left(y\right)}&{\theta\xi_{1}\left(y\right)}\\ {\theta\xi_{2}\left(y\right)}&{\alpha_{2}\left(y\right)+\theta^{2}n_{2}\left(y\right)}\end{array}\right)$ (1.3)
is a chiral “twisted superfield” whose component fields are U(3) matrix functions (e.g. $\alpha_{1}=\alpha_{1}^{A}t^{A}$) of $y^{\mu}=x^{\mu}+i\theta\sigma^{\mu}\bar{\theta}$. The twisted supergauge transformation of eq (1.2) maintains the boson-fermion structure of the real twisted superfield as well as its group structure. To the latter point, if the group was SU(3)$\times$×SU(3) instead of U(3)$\times$×U(3), the supergauge transformation would not be consistent, since a general supergauge transformation would generate terms in each block proportional to $t^{0}$. On the other hand, a U(3)$\times$×U(3) twisted supergauge transformation is consistent.
Like normal real superfields, the real twisted superfield supports conjugate representations. To see this, it is helpful to follow the presentation of [8] and re-express an infinitesimal twisted supergauge transformation as
$V$ $\to V+i\Lambda^{{\dagger}}-i\Lambda-{\tfrac{1}{2}}i\left[V,\left(\Lambda^{{\dagger}}+\Lambda\right)\right]$ (1.4)
$+i\sum_{k=1}^{\infty}\frac{B_{2k}}{\left(2k\right)!}\left[V,\left[V,...\left[V,\left(\Lambda^{{\dagger}}-\Lambda\right)\right]...\right]\right],$
where $B_{2k}$ are Bernoulli numbers. Both $V$ and $\Lambda$ can be expanded in terms of component fields multiplied by U(6) matrices $T^{X}$, where the index X runs over the 36 adjoint indices of U(6). As with any unitary group, the same structure functions $f^{XYZ}$ satisfy both $\left[T^{X},T^{Y}\right]=if^{XYZ}T^{Z}$ and $\left[-T^{XT},-T^{YT}\right]=-if^{XYZ}T^{ZT}$. Since products of matrices in eq (1.4) only enter by way of commutators, a conjugate representation is available by replacing each $T^{X}$ in eq (1.4) with $\left(-T^{X}\right)^{T}$. In other words, a twisted real superfield that transforms by eq (1.2) also transforms as follows:
$e^{-V}\to e^{i\Lambda}e^{-V}e^{-i\Lambda^{{\dagger}}}.$ (1.5)
Despite the fact that (1.1) and (1.3) are not N=1 superfields and (1.2) is not a normal supergauge transformation, the word “twisted” will be dropped for brevity in much of the rest of the paper.
A consequence of eq (1.4) is the fact that one component of the real superfield has a supergauge transformation independent of the other components. Taking the trace of eq (1.4), one finds
${\tfrac{1}{\sqrt{3}}}{\rm Tr}\left(V\right)=V_{+}^{0}\to V_{+}^{0}+i\Lambda_{+}^{0{\dagger}}-i\Lambda_{+}^{0}$ (1.6)
$\Lambda_{+}^{0}={\tfrac{1}{\sqrt{3}}}{\rm Tr}\left(\Lambda\right),$
where ${\rm Tr}$ is the 6$\times$6 trace. The reason that this supergauge transformation is independent is because all of the commutators in eq (1.4) are proportional to some 6$\times$6 traceless matrix, so none of them can contribute to eq (1.6). Since by definition, the fermions of the real superfield are all in off-diagonal blocks, the field $V_{+}^{0}$ does not include any fermions, only bosons. Inside of $V$, the field $V_{+}^{0}$ is multiplied the 6$\times$6 matrix $T_{+}^{0}$ defined via
$T_{\pm}^{A}={\tfrac{1}{\sqrt{2}}}\left(T_{1}^{A}\pm T_{2}^{A}\right)$ (1.7)
$T_{1}^{A}=\left(\begin{array}[]{cc}{t^{A}}&{0}\\ {0}&{0}\end{array}\right)\,\,\,\,T_{2}^{A}=\left(\begin{array}[]{cc}{0}&{0}\\ {0}&{t^{A}}\end{array}\right).$
This definition provides another way of saying that $T_{+}^{0}$ is ${\tfrac{1}{2\sqrt{3}}}$ of the 6$\times$6 unit matrix.
As is often done in superspace gauge theories, the real superfield will be rescaled to explicitly show the coupling constant. In this case, the following rescaling is performed:
$V\to 2gV^{\prime}+2g_{+}V_{+}^{0}T_{+}^{0}$ (1.8)
$V^{\prime}=V-V_{+}^{0}T_{+}^{0}.$
Since $V_{+}^{0}$ has its own, independent supergauge transformation, it also has its own coupling constant.
Now that the gauge transformation properties of the real superfield have been identified, gauge invariant action terms can be defined. Just as with normal superfields, the following chiral twisted superfields can be defined:
$W^{\prime}_{\alpha}$ $=-{\tfrac{1}{8g}}i\bar{D}^{2}\left(e^{-2gV^{\prime}}D_{\alpha}e^{2gV^{\prime}}\right)$ (1.9)
$W_{\alpha}^{0}$ $=-{\tfrac{1}{4}}i\bar{D}^{2}D_{\alpha}V_{+}^{0},$
where $D_{\alpha}=\partial_{\alpha}+i\sigma_{\alpha\dot{\alpha}}^{\mu}\bar{\theta}^{\dot{\alpha}}\partial_{\mu}$. Under a supergauge transformation, these fields transform as follows:
$W^{\prime}_{\alpha}$ $\to e^{i\Lambda}W^{\prime}_{\alpha}e^{-i\Lambda}$ (1.10)
$W_{\alpha}^{0}$ $\to W_{\alpha}^{0}.$
As a result, the following terms in the action are supergauge invariant:
$S_{V}=$ $-{\tfrac{1}{2}}\int d^{4}xd^{2}\theta\left(1+4m_{\lambda}\theta^{2}\right){\rm Tr}\left(W^{\prime\alpha}W^{\prime}_{\alpha}\right)$ (1.11)
$-{\tfrac{1}{4}}\int d^{4}xd^{2}\theta W^{0\alpha}W_{\alpha}^{0}+h.c.,$
where $h.c.$ stands for Hermitian conjugate and $m_{\lambda}$ is a “gaugino mass”.
The action may also include the following gauge-invariant Fayet Iliopoulos term:
$S_{\xi}=\tfrac{1}{\sqrt{3}}\xi_{+}\int d^{4}xd^{2}\theta d^{2}\bar{\theta}g_{+}V_{+}^{0}.$ (1.12)
In addition to the real superfield, the theory includes the following chiral twisted superfield in an adjoint representation of twisted U(3)$\times$×U(3):
$\Phi=\left(\begin{array}[]{cc}{\varphi_{1}\left(y\right)+\theta^{2}f_{1}\left(y\right)}&{-\sqrt{2}i\theta\chi\left(y\right)}\\ {-\sqrt{2}i\theta\tilde{\chi}\left(y\right)}&{\varphi_{2}\left(y\right)+\theta^{2}f_{2}\left(y\right)}\end{array}\right)$ (1.13)
The adjoint superfield transforms as follows:
$\Phi\to e^{i\Lambda}\Phi e^{-i\Lambda}.$ (1.14)
The following action terms involving this field are supergauge invariant:
$S_{\Phi}$ $=2\int d^{4}xd^{2}\theta d^{2}\bar{\theta}\times$ (1.15)
×$\times{\rm Tr}\Big{(}\Phi^{{\dagger}}e^{2gV^{\prime}}\Phi e^{-2gV^{\prime}}(1-\theta^{2}\bar{\theta}^{2}\sum_{m}m^{2}_{\Phi m}\sqrt{\tfrac{3}{2}}T_{m}^{0})\Big{)}$
$-2\int d^{4}xd^{2}\theta{\rm Tr}\left({\tfrac{1}{2}}m_{\Phi}\Phi^{2}+{\tfrac{1}{3}}\Gamma_{\Phi}\Phi^{3}\right)+h.c.$
The $m^{2}_{\Phi m}$ terms are scalar mass terms. Despite the explicit group matrices $T_{m}^{0}$, the terms are supergauge invariant since the factor of $\theta^{2}\bar{\theta}^{2}$ limits gauge transformations to ones that remain within the same 3$\times$3 diagonal block. Just like the gaugino mass term (and like analogous soft supersymmetry breaking terms), the scalar mass terms break superspace translation invariance (which is not imposed in this model anyway), but do not break superspace gauge invariance.
In addition to adjoint-representation chiral fields, the theory also includes three flavors of 6-vector chiral twisted superfields in the fundamental and anti-fundamental representations of twisted U(3)$\times$×U(3):
$Q_{1F}=\left(\begin{array}[]{c}{\phi_{1F}+\theta^{2}f_{1F}}\\ {\sqrt{2}\theta\psi_{2F}}\end{array}\right)$ (1.16)
$Q_{2F}=\left(\begin{array}[]{c}{\sqrt{2}\theta\psi_{1F}}\\ {\phi_{2F}+\theta^{2}f_{2F}}\end{array}\right)$
$\tilde{Q}_{1F}=\left(\begin{array}[]{cc}{\tilde{\phi}_{1F}+\theta^{2}\tilde{f}_{1F},}&{\sqrt{2}\theta\tilde{\psi}_{2F}}\end{array}\right)$
$\tilde{Q}_{2F}=\left(\begin{array}[]{cc}{\sqrt{2}\theta\tilde{\psi}_{1F},}&{\tilde{\phi}_{2F}+\theta^{2}\tilde{f}_{2F}}\end{array}\right)$
where $F\in\left\{1,2,3\right\}$ is a flavor index, and each component field is a chiral 3-vector (or covector). The supergauge transformation for each of these superfields depends upon their flavor in the following way:
$\tilde{Q}_{mF}\to\tilde{Q}_{mF}e^{-i\left(\Lambda^{\prime}+q_{F}\Lambda_{+}^{0}T_{+}^{0}\right)}$ (1.17)
$Q_{mF}\to e^{i\left(\Lambda^{\prime}+q_{F}\Lambda_{+}^{0}T_{+}^{0}\right)}Q_{mF}$
$q_{1}=-2,\,\,\,\,q_{2}=q_{3}=1,$
where $m\in\left\{1,2\right\}$. As a result, the following terms in the action are supergauge invariant:
$S_{Q}$ $=\sum_{mF}\int d^{4}xd^{2}\theta d^{2}\bar{\theta}\times$ (1.18)
×$\hskip 11.38092pt\times\Big{(}\left(1-m^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)Q_{mF}^{{\dagger}}e^{2\left(gV^{\prime}+q_{F}g_{+}V_{+}^{0}T_{+}^{0}\right)}Q_{mF}$
$\hskip 14.22636pt+\left(1-\tilde{m}^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)\tilde{Q}_{mF}e^{-2\left(gV^{\prime}+q_{F}g_{+}V_{+}^{0}T_{+}^{0}\right)}\tilde{Q}_{mF}^{{\dagger}}\Big{)}$
$-\sum_{m}\int d^{4}xd^{2}\theta\Bigg{(}\tilde{Q}_{m1}\left(m_{m11}+\sqrt{2}\Gamma_{m11}\Phi\right)Q_{m1}$
$+\sum_{F,F^{\prime}=2,3}\tilde{Q}_{mF}\left(m_{mFF^{\prime}}+\sqrt{2}\Gamma_{mFF^{\prime}}\Phi\right)Q_{mF^{\prime}}\Bigg{)}$
$+h.c.$
To maintain gauge invariance in light of the transformations in eq (1.17), the above superpotential terms do not connect flavor 1 fields with flavor 2 or 3 fields. The $m^{2}_{mF}$ and $\tilde{m}^{2}_{mF}$ terms generate additional mass terms for the fundamental and conjugate scalars.
The theory presented above is free of gauge anomalies. There is a simple reason: for every fermion in the theory, there is another fermion in a conjugate representation with opposite Abelian charges. Since the theory is a gauge theory and is free of gauge anomalies, it is assumed that the theory is renormalizable.
Since all of the action terms presented above are invariant to twisted supergauge transformations, it is possible to restrict the real superfield to a Wess-Zumino-like gauge. In that gauge, the real superfield takes the form:
$V=\left(\begin{array}[]{cc}{-\bar{\theta}\bar{\sigma}^{\mu}A_{1\mu}\theta+{\tfrac{1}{2}}\bar{\theta}^{2}d_{1}\theta^{2}}&{i\bar{\theta}\tilde{\lambda}^{{\dagger}}\theta^{2}-i\bar{\theta}^{2}\lambda\theta}\\ {i\bar{\theta}\lambda^{{\dagger}}\theta^{2}-i\bar{\theta}^{2}\tilde{\lambda}\theta}&{-\bar{\theta}\bar{\sigma}^{\mu}A_{2\mu}\theta+{\tfrac{1}{2}}\bar{\theta}^{2}d_{2}\theta^{2}}\end{array}\right).$ (1.19)
In [4], it was shown that a Wess-Zumino-like gauge is accessible for a theory with this kind of twisted supergauge invariance. After imposition of this Wess-Zumino gauge, the residual gauge invariance is just local spacetime gauge invariance. In the following, the fermions $\tilde{\lambda}$ and $\lambda$ will be referred to as “gauginos” despite the fact that they are in the (3,3*) representation of the gauge group, rather than the adjoint representation.
Although the Abelian field $V_{+}^{0}$ has its own, independent gauge transformation, the second Abelian field $V_{-}^{0}$ (group structure $T_{-}^{0}$) does not decouple from the gauge transformations described above. That is why the same designation $g$ was used for both the $V_{-}^{0}$ and nonAbelian couplings above. Nonetheless, in the Wess-Zumino gauge, since $V_{-}^{0}$ is an Abelian field, it can accommodate different charges multiplying its coupling constant when acting on different chiral fields. These charges (labelled as $q^{\prime}_{F}$ and $q^{\prime}_{A}$) are chosen to be:
$q^{\prime}_{1}=2,\,\,\,q^{\prime}_{2}=1,\,\,\,q^{\prime}_{3}=-1,\,\,\,q^{\prime}_{A}=-6\,,$ (1.20)
where $q^{\prime}_{A}$ is the charge multiplying $V_{-}^{0}$ when interacting with chiral adjoint fields $\Phi$. The reason for these choices is to reproduce the X17 data discussed in section 5. If instead, all charges for $V_{-}^{0}$ are chosen to be $1$, the theory is not able to reproduce all of the X17 data, but almost all of the other results of the paper still hold.
The classical theory has 48 parameters that can be adjusted classically: 2 coupling constants, 8 Abelian charges, 26 masses, 11 superpotential couplings, and a Fayet-Iliopoulos term. In the unification section of the paper, it is argued that both gauge couplings may be the same at the unification scale ($g_{+}=g$). It is also assumed that the following 5 parameters are zero classically:
$m_{23F}=m_{223}=m_{23}=\tilde{m}_{23}=0$ (1.21)
It is possible that many of the remaining parameters are also zero classically but are generated quantum mechanically.
2 Dynamical Symmetry Breaking
This section identifies a minimum of the scalar potential that breaks the gauge symmetry in stages from SU(3)$\times$×SU(3)$\times$×U(1)$\times$×U(1) down to SU(3)$\times$×U(1) and labels the fermions in the model based on their Standard Model symmetries.
The scalar potential for this model can be expressed in terms of its auxiliary fields and scalar mass terms:
$\mathcal{V}$ $={\tfrac{1}{2}}\left(d_{+}^{0}\right)^{2}+{\tfrac{1}{2}}\left(d_{-}^{0}\right)^{2}+{\tfrac{1}{2}}\sum_{m,a}\left(d_{m}^{a}\right)^{2}$ (2.1)
$+\sum_{mF}\left(f_{mF}^{{\dagger}}f_{mF}+\tilde{f}_{mF}\tilde{f}_{mF}^{{\dagger}}\right)+2{\rm tr}\left(f_{1}^{{\dagger}}f_{1}\right)+2{\rm tr}\left(f_{2}^{{\dagger}}f_{2}\right)$
$+\sum_{mF}\left(m^{2}_{mF}\phi^{\dagger}_{mF}\phi_{mF}+\tilde{m}^{2}_{mF}\tilde{\phi}^{\dagger}_{mF}\tilde{\phi}_{mF}\right)$
$+m^{2}_{\Phi 1}{\rm tr}(\varphi_{1}^{\dagger}\varphi_{1})+m^{2}_{\Phi 2}{\rm tr}(\varphi_{2}^{\dagger}\varphi_{2}),$
where $d_{\pm}^{0}={\tfrac{1}{\sqrt{2}}}\left(d_{1}^{0}\pm d_{2}^{0}\right)$, and lower-case ${\rm tr}$ defines a 3$\times$3 trace. By their equations of motion, the auxiliary fields are equal to linear or quadratic functions of the scalar fields. For example, the equations of motion for the d terms of $\mathcal{V}$ result in:
$-d_{+}^{0}$ $={\frac{g_{+}}{\sqrt{12}}}\left(\xi_{+}+\sum_{mF}q_{F}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)\right)$ (2.2)
$-d_{-}^{0}$ $=-{\frac{g}{\sqrt{12}}}\sum_{mF}q^{\prime}_{F}\left(-1\right)^{m}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)$
$-d_{m}^{a}$ $=2g{\rm tr}\left(t^{a}\left[\varphi_{m},\varphi_{m}^{{\dagger}}\right]\right)$
$+g\sum_{F}\left(\phi_{mF}^{{\dagger}}t^{a}\phi_{mF}-\tilde{\phi}_{mF}t^{a}\tilde{\phi}_{mF}^{{\dagger}}\right)\,.$
The Abelian charges $q_{F}$ and $q^{\prime}_{F}$ for $V_{\pm}^{0}$ (including $d_{\pm}^{0}$) were defined in eqs (1.17) and (1.20). Just as for the $d$ auxiliary fields, the equations of motion can also be used to derive expressions for the $f$ auxiliary fields in terms of scalar fields.
It is assumed that the masses $m_{1F}$ and $\tilde{m}_{1F}$ are large compared to $\xi_{+}$. In that case, the minimum of the scalar potential is achieved when the fundamental and conjugate scalars with an $m=1$ subscript have no vacuum expection value (vev):
$\left\langle\tilde{\phi}_{1F}\right\rangle=\left\langle\phi_{1F}\right\rangle=0,$ (2.3)
where $\left\langle\phi_{mF}\right\rangle$ denotes the vev of $\phi_{mF}$.
Due to eq (1.21) along with the assumption that $m_{21}$ and $\tilde{m}_{22}$ are small, it is assumed that the $m=2$ fundamental and conjugate scalars acquire vevs. Following precedent from Supersymmetric QCD (SQCD) [14, 13], the vevs in the $m=2$ sector are assumed to take the following form:
$\left\langle\tilde{\phi}_{2F}\right\rangle^{n}$ $=i\delta_{F}^{n}\bar{\tilde{\phi}}_{2F}$ (2.4)
$\left\langle\phi_{2F}\right\rangle_{n}$ $=-i\delta_{Fn}\bar{\phi}_{2F},$
where an overbar on a component of a scalar field (e.g. $\bar{\phi}_{2F}$ above) is used to denote the magnitude (real, positive) of the vev of that component. In eq (2.4), the index $n$ represents the SU(3) index of the 3-vectors $\tilde{\phi}_{2F}$ and $\phi_{2F}$. For example, writing out the SU(3) “color” components: $\left\langle\tilde{\phi}_{23}\right\rangle=i\left(0,0,\bar{\tilde{\phi}}_{23}\right)$. In other words, the vevs $\left\langle\tilde{\phi}_{2F}\right\rangle^{n}$ and $\left\langle\phi_{2F}\right\rangle_{n}$ form 3$\times$3 diagonal matrices in their flavor-“color” indices. The word “color” is being used here in order to make a connection with SQCD techniques, but in this model after symmetry breaking, the 3 “color” indices of the $m=2$ scalars will actually correspond to 2 isospin doublet indices and 1 singlet index. The phases of the vevs are chosen to simplify fermion mass matrices in the next section.
Although $m_{21}$ and $\tilde{m}_{22}$ are assumed to be small, it is assumed that the scalar masses $\tilde{m}_{21}$ and $m_{22}$ are large compared to $\xi_{+}$. To accommodate a nontrivial minimum in the presence of these masses, the following is assumed classically:
$\bar{\tilde{\phi}}_{21}=\bar{\phi}_{22}=0\hskip 14.22636pt\textrm{classically}.$ (2.5)
In the appendix, it is argued that small vevs are generated for these fields quantum mechanically.
The following adjoint vevs are considered for the classical theory:
$\left\langle\varphi_{1}\right\rangle$ $=0\hskip 14.22636pt\textrm{classically}$ (2.6)
$\left\langle\varphi_{2}\right\rangle$ $=\frac{i\bar{\varphi}_{2}}{\sqrt{2}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{0}\\ {0}&{1}&{0}\end{array}\right)\,.$
In the next section, in light of the quantum vacuum moduli space, different assumptions will be made for both $\left\langle\varphi_{1}\right\rangle$ and $\left\langle\varphi_{2}\right\rangle$ that have ${\rm tr}\left\langle\varphi_{m}^{2}\right\rangle\neq 0$.
Assuming $m_{211}$, $m_{222}$ and $m_{\Phi}$ are small classically, the vev of the classical scalar potential is equal to
$2\left\langle\mathcal{V}\right\rangle$ $={\tfrac{1}{12}}g_{+}^{2}\left(\xi_{+}-2\bar{\phi}_{21}^{2}-\bar{\tilde{\phi}}_{22}^{2}+\Delta\bar{\phi}_{3}^{2}\right)^{2}$ (2.7)
$+{\tfrac{1}{6}}g^{2}\left(2\bar{\phi}_{21}^{2}-\Delta\bar{\phi}_{3}^{2}\right)^{2}+{\tfrac{1}{6}}g^{2}\bar{\tilde{\phi}}_{22}^{4}$
$+{\tfrac{1}{4}}g^{2}\left(\bar{\tilde{\phi}}_{22}^{2}+\Delta\bar{\phi}_{3}^{2}+\bar{\varphi}_{2}^{2}\right)^{2}$
$+2m_{21}^{2}\bar{\phi}_{21}^{2}+2\tilde{m}_{22}^{2}\bar{\tilde{\phi}}_{22}^{2}+m_{\Phi 2}^{2}\bar{\varphi}_{2}^{2}$
$\Delta\bar{\phi}_{3}^{2}$ $=\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}\,.$
The minimum of the above classical scalar potential at the unification scale is found when the parameters and scalar vevs satisfy the following equations:
$m^{2}_{\Phi 2}=$ $m^{2}_{21}$ (2.8)
$\tilde{m}_{22}^{2}=$ $\tfrac{1}{3}g^{2}\bar{\phi}^{2}_{21}+\tfrac{1}{6}g^{2}\bar{\varphi}_{2}^{2}-\tfrac{2}{3}m^{2}_{\Phi 2}$
$\tfrac{1}{3}\xi_{+}=$ $2\bar{\phi}^{2}_{21}+\tfrac{4}{3}\bar{\tilde{\phi}}_{22}^{2}+\bar{\varphi}_{2}^{2}$
$\Delta\bar{\phi}_{3}^{2}=$ $-\bar{\tilde{\phi}}_{22}^{2}-\bar{\varphi}_{2}^{2}-2m^{2}_{\Phi 2}/g^{2}\,,$
where based on the unification section, it has been assumed that $g_{+}=g$.
Theoretically, the vacuum chooses the values of all of the scalar vevs, subject to the restrictions of eq (2.8). Phenomenologically, the goal is to determine vacuum choices that would result in the observed particle spectrum and symmetries. From that point of view, all vevs and parameters except those on the left side of eq (2.8) may be chosen independently. The equations then determine the parameters on the left of eq (2.8). For the present model, three of the vevs are chosen to be at the following scales:
$\begin{array}[]{|r|c|l|}\hline\cr\rm{vev}&\textrm{vev scale}&\mbox{Symmetry Breaking}\\ \hline\cr g\bar{\tilde{\phi}}_{23}&\sim 10^{15}\,\rm{GeV}&\mbox{Unification: SU(3)$\times$SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)}\\ g_{2}\bar{\phi}_{21}&\sim 10^{2}\,\rm{GeV}&\mbox{Electroweak: SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)$\times$U(1)}\\ g_{Z^{\prime}}\bar{\tilde{\phi}}_{22}&\sim 10^{-2}\,\rm{GeV}&\mbox{$Z^{\prime}$ acquires mass: SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)}\\ \hline\cr\end{array}$ (2.9)
where $g_{2}$ is the Weak coupling and $g_{Z^{\prime}}$ is the $Z^{\prime}$ coupling that is driven to a very small value by a heavy right-handed neutrino, as discussed in section 4. The last equation in eq (2.8) causes the following vev to also be at the unification scale:
$g\bar{\phi}_{23}\sim 10^{15}\rm{GeV}.$ (2.10)
In Supersymmetric QCD (SQCD) with fewer flavors than colors, the classical vacuum does not determine actual values of vevs, but just determines differences like $\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$ in eq (2.8). But the SQCD quantum vacuum causes the vevs involved in those differences to get very large. In fact, the quantum vacuum drives them to infinity, so that for fewer flavors than colors, SQCD does not have a vacuum [14, 13]. In the appendix, it is pointed out that similar forces are at work in this theory, but small counterbalancing forces are also at work. The net result is that the quantum vacuum requires the vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ to become very large, but not infinite. This is justification for why these two fields have vevs at a very large unification scale.
Section 4 of this paper determines the numerical value of the unification scale by starting at electroweak energies and running the SU(2) and SU(3) coupling constants up to the scale where they become the same. In that section, it is argued that the U(1) coupling $g_{+}$ may also unify with the nonAbelian couplings at that same scale.
A difference from the Standard Model in the above symmetry breaking is that there is an extra U(1) field (the $Z^{\prime}$) that acquires a mass well below the electroweak scale. This will be discussed in more detail later in this section and in section 5.
In the above symmetry breaking, the SU(3) gluons of the Standard Model come from the $A^{a}_{1\mu}$ gauge bosons, while the SU(2) weak fields come from the $A^{a}_{2\mu}$ gauge bosons. The U(1) fields in the model are a mixture of $A^{0}_{1\mu}$ and $A^{A}_{2\mu}$ gauge bosons. The progression of the U(1) fields through the various stages of symmetry breaking is discussed in detail below.
Via the Brout-Englert-Higgs mechanism, the scalar vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ impart unification-scale masses to $A_{2\mu}^{4},A_{2\mu}^{5},A_{2\mu}^{6}$ and $A_{2\mu}^{7}$ as well as to one diagonal gauge boson. Consequently, the gauge symmetry is broken down to SU(3)$\times$×SU(2)$\times$×U(1)$\times$×U(1). To see the group structure of the remaining massless diagonal gauge fields, it is helpful to use the notation of eq (1.7) and re-expand the gauge fields into the linear combinations below:
$\left(\begin{array}[]{c}{A_{\mu}^{Y}}\\ {A_{\mu}^{Y^{\prime}}}\\ {A_{\mu}^{U}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{U}}&{\sin\phi_{U}}\\ {0}&{-\sin\phi_{U}}&{\cos\phi_{U}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{U}}&{0}&{\sin\theta_{U}}\\ {0}&{1}&{0}\\ {-\sin\theta_{U}}&{0}&{\cos\theta_{U}}\end{array}\right)\left(\begin{array}[]{c}{A_{+\mu}^{0}}\\ {A_{-\mu}^{0}}\\ {A_{2\mu}^{8}}\end{array}\right),$ (2.11)
where $A_{\mu}^{U}$ acquires a unification-scale mass, but $A_{\mu}^{Y}$ and $A_{\mu}^{Y^{\prime}}$ remain massless. In order to achieve the relation
$q_{F}g_{+}A_{+\mu}^{0}T_{+}^{0}+q^{\prime}_{F}gA_{-\mu}^{0}T_{-}^{0}+gA_{2\mu}^{8}T_{2}^{8}=g_{Y}A_{\mu}^{Y}T_{F}^{Y}+g_{Y^{\prime}}A_{\mu}^{Y^{\prime}}T_{F}^{Y^{\prime}}+g_{U}A_{\mu}^{U}T_{F}^{U},$ (2.12)
appearing in the action, the coupling constants and group matrices must satisfy:
$\left(\begin{array}[]{c}{g_{Y}T_{F}^{Y}}\\ {g_{Y^{\prime}}T_{F}^{Y^{\prime}}}\\ {g_{U}T_{F}^{U}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{U}}&{\sin\phi_{U}}\\ {0}&{-\sin\phi_{U}}&{\cos\phi_{U}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{U}}&{0}&{\sin\theta_{U}}\\ {0}&{1}&{0}\\ {-\sin\theta_{U}}&{0}&{\cos\theta_{U}}\end{array}\right)\left(\begin{array}[]{c}{q_{F}g_{+}T_{+}^{0}}\\ {q^{\prime}_{F}gT_{-}^{0}}\\ {gT_{2}^{8}}\end{array}\right).$ (2.13)
This is just a generalization of a Weinberg angle rotation. A more complete generalization could involve a third angle mixing the two massless fields, but that is not needed here. The reason that the group matrices on the left have a flavor dependence is because eq (2.12) involves the flavor-dependent Abelian charges $q_{F}$ and $q^{\prime}_{F}$.
The angles $\theta_{U}$ and $\phi_{U}$ in eq (2.13) are chosen so that $T^{Y}$ and $T^{Y^{\prime}}$ (for flavor 3) have zeros in their sixth diagonal slot, so that they get no mass contribution from $\bar{\phi}_{23}$ or $\bar{\tilde{\phi}}_{23}$. Specifically, the angles are given by:
$\tan\theta_{U}$ $=g_{+}/2g=\tfrac{1}{2}$ (2.14)
$\tan\phi_{U}$ $=\tfrac{1}{2}\cos\theta_{U}=\tfrac{1}{\sqrt{5}}.$
where the second equalities above assume that $g_{+}=g$ at the unification scale. In that case $g_{Y}=g_{Y^{\prime}}=g$, and the group matrices take the forms:
$T^{Y}$ $={\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}{\rm diag}\left({\tfrac{2}{3}},{\tfrac{2}{3}},{\tfrac{2}{3}},1,1,0\right)$ (2.15)
$T_{1}^{Y}$ $=-{\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}{\rm diag}\left({\tfrac{4}{3}},{\tfrac{4}{3}},{\tfrac{4}{3}},1,1,2\right)$
$T_{3}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(-1,-1,-1,1,1,0\right)$
$T_{1}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(2,2,2,-1,-1,-2\right)$
$T^{Y}$ is the $A_{\mu}^{Y}$ group structure when acting on gaugino, adjoint, flavor 2 or flavor 3 fields. $T_{1}^{Y}$ is the $A_{\mu}^{Y}$ group structure when acting on flavor 1 fields. $T_{3}^{Y^{\prime}}$ and $T_{1}^{Y^{\prime}}$ are the $A_{\mu}^{Y^{\prime}}$ group structures when acting on flavor 3 and 1 fields. The $A_{\mu}^{Y^{\prime}}$ group structures when acting on gauginos, adjoint, or flavor 2 chiral fields can be calculated from eqs (1.20), (2.13) and (2.14).
It will be seen below that the $T^{Y}$ and $T_{1}^{Y}$ matrices have the correct form for their gauge boson $A_{\mu}^{Y}$ to be identified as the U(1) weak hypercharge field of the Standard Model with $-{\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}g_{Y}$ identified as the weak hypercharge coupling. The $A_{\mu}^{Y^{\prime}}$ gauge field with its coupling $g_{Y^{\prime}}$ is a second U(1) gauge boson in this model that remains massless at the unification scale. Due to eq (2.11), there is no mixing between the $A_{\mu}^{Y}$ and $A_{\mu}^{Y^{\prime}}$ gauge fields. Below the unification scale, the couplings for the SU(3), SU(2), $Y$ and $Y^{\prime}$ groups run differently, so they are denoted by $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$.
In this model according to eq (2.9), electroweak symmetry is primarily broken by $\bar{\phi}_{21}$. This vev gives masses to the $W$ and $Z$ bosons, leaving only the SU(3) gluons, the photon and the $Z^{\prime}$ boson massless. To see the structure of the diagonal fields, one may again make a Weinberg-angle rotation:
$\left(\begin{array}[]{c}{eT_{F}^{\gamma}}\\ {g_{Z^{\prime}}T_{F}^{Z^{\prime}}}\\ {g_{Z}T_{F}^{Z}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{Z}}&{\sin\phi_{Z}}\\ {0}&{-\sin\phi_{Z}}&{\cos\phi_{Z}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{Z}}&{0}&{\sin\theta_{Z}}\\ {0}&{1}&{0}\\ {-\sin\theta_{Z}}&{0}&{\cos\theta_{Z}}\end{array}\right)\left(\begin{array}[]{c}{g_{Y}T_{F}^{Y}}\\ {g_{Y^{\prime}}T_{F}^{Y^{\prime}}}\\ {g_{2}T_{2}^{3}}\end{array}\right).$ (2.16)
The angle $\theta_{Z}$ is chosen to make $(T_{1}^{\gamma})_{44}=(T^{\gamma})_{55}=0$, so that the photon gets no mass from $\bar{\phi}_{21}$, $\bar{\tilde{\phi}}_{22}$ or $\bar{\varphi}_{2}$.
The angle $\phi_{Z}$ is chosen to make $(T_{1}^{Z^{\prime}})_{44}=0$ and $(T_{A}^{Z^{\prime}})_{55}=(T_{A}^{Z^{\prime}})_{66}$, so that the $Z^{\prime}$ gets no mass from $\bar{\phi}_{21}$ or $\bar{\varphi}_{2}$. The resulting angles are given by:
$\tan\theta_{Z}$ $={\sqrt{{\tfrac{3}{5}}}g_{Y}/g_{2}}$ (2.17)
$\tan\phi_{Z}$ $=\tfrac{2}{\sqrt{10}}\cos\theta_{Z}g_{Y^{\prime}}/g_{2}.$
In section 4, it will be argued that $g_{Y^{\prime}}$ will be driven to a very small value from an effective anomaly generated by a very massive right-handed neutrino. As a result, the angle $\phi_{Z}$ is very small, and $\theta_{Z}$ is very close to the Weinberg angle $\theta_{W}$ of the Standard Model.
With the following coupling constant normalization,
$e=-g_{2}\sin\theta_{Z}\,,$ (2.18)
the photon group structure is given by
$eT^{\gamma}$ $=-e{\rm diag}\left({\tfrac{1}{3}},{\tfrac{1}{3}},{\tfrac{1}{3}},1,0,0\right)$ (2.19)
$eT_{1}^{\gamma}$ $=e{\rm diag}\left({\tfrac{2}{3}},{\tfrac{2}{3}},{\tfrac{2}{3}},0,1,1\right)\,.$
For the case where $|\phi_{Z}|\ll 1,$ the Z boson group structure is approximately:
$g_{Z}T^{Z}$ $\simeq-(g_{2}/\cos\theta_{W}){\rm diag}\left({\tfrac{1}{3}x},{\tfrac{1}{3}x},{\tfrac{1}{3}x},-\tfrac{1}{2}+x,\tfrac{1}{2},0\right)$ (2.20)
$g_{Z}T_{1}^{Z}$ $\simeq(g_{2}/\cos\theta_{W}){\rm diag}\left({\tfrac{2}{3}x},{\tfrac{2}{3}x},{\tfrac{2}{3}x},\tfrac{1}{2},-\tfrac{1}{2}+x,x\right)$
$|\phi_{Z}|$ $\ll 1,$
where $x=\sin^{2}\theta_{Z}\simeq\sin^{2}\theta_{W}$. The group structure of the $Z^{\prime}$ boson is shown in eq (5.1) of section 5.
[…]
$\tilde{\lambda}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}{\tilde{u}_{W1}^{G}}&{\tilde{u}_{W2}^{G}}&{\tilde{u}_{W3}^{G}}\\ {\tilde{d}_{W1}^{G}}&{\tilde{d}_{W2}^{G}}&{\tilde{d}_{W3}^{G}}\\ {\tilde{d}_{1}^{G}}&{\tilde{d}_{2}^{G}}&{\tilde{d}_{3}^{G}}\end{array}\right)\,\,\,\,\lambda=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}{u_{W1}^{G}}&{d_{W1}^{G}}&{d_{1}^{G}}\\ {u_{W2}^{G}}&{d_{W2}^{G}}&{d_{2}^{G}}\\ {u_{W3}^{G}}&{d_{W3}^{G}}&{d_{3}^{G}}\end{array}\right)$ (2.21)
$\tilde{\chi}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}{\tilde{u}_{W1}^{A}}&{\tilde{u}_{W2}^{A}}&{\tilde{u}_{W3}^{A}}\\ {\tilde{d}_{W1}^{A}}&{\tilde{d}_{W2}^{A}}&{\tilde{d}_{W3}^{A}}\\ {\tilde{d}_{1}^{A}}&{\tilde{d}_{2}^{A}}&{\tilde{d}_{3}^{A}}\end{array}\right)\,\,\,\,\chi=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}{u_{W1}^{A}}&{d_{W1}^{A}}&{d_{1}^{A}}\\ {u_{W2}^{A}}&{d_{W2}^{A}}&{d_{2}^{A}}\\ {u_{W3}^{A}}&{d_{W3}^{A}}&{d_{3}^{A}}\end{array}\right)$
$\psi_{11}=\left(\begin{array}[]{c}{u_{1}^{\left(1\right)}}\\ {u_{2}^{\left(1\right)}}\\ {u_{3}^{\left(1\right)}}\end{array}\right)\,\psi_{12}=\left(\begin{array}[]{c}{d_{1}^{\left(2\right)}}\\ {d_{2}^{\left(2\right)}}\\ {d_{3}^{\left(2\right)}}\end{array}\right)\,\psi_{13}=\left(\begin{array}[]{c}{d_{1}^{\left(3\right)}}\\ {d_{2}^{\left(3\right)}}\\ {d_{3}^{\left(3\right)}}\end{array}\right)\,\psi_{21}=\left(\begin{array}[]{c}{\nu_{W}^{\left(1\right)}}\\ {e_{W}^{\left(1\right)}}\\ {e^{\left(1\right)}}\end{array}\right)\,\psi_{22}=\left(\begin{array}[]{c}{e_{W}^{\left(2\right)}}\\ {\nu_{W}^{\left(2\right)}}\\ {\nu^{\left(2\right)}}\end{array}\right)\,\psi_{23}=\left(\begin{array}[]{c}{e_{W}^{\left(3\right)}}\\ {\nu_{W}^{\left(3\right)}}\\ {\nu^{\left(3\right)}}\end{array}\right)$
$\tilde{\psi}_{11}^{T}=\left(\begin{array}[]{c}{\tilde{u}_{1}^{\left(1\right)}}\\ {\tilde{u}_{2}^{\left(1\right)}}\\ {\tilde{u}_{3}^{\left(1\right)}}\end{array}\right)\,\tilde{\psi}_{12}^{T}=\left(\begin{array}[]{c}{\tilde{d}_{1}^{\left(2\right)}}\\ {\tilde{d}_{2}^{\left(2\right)}}\\ {\tilde{d}_{3}^{\left(2\right)}}\end{array}\right)\,\tilde{\psi}_{13}^{T}=\left(\begin{array}[]{c}{\tilde{d}_{1}^{\left(3\right)}}\\ {\tilde{d}_{2}^{\left(3\right)}}\\ {\tilde{d}_{3}^{\left(3\right)}}\end{array}\right)\,\tilde{\psi}_{21}^{T}=\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{\left(1\right)}}\\ {\tilde{e}_{W}^{\left(1\right)}}\\ {\tilde{e}^{\left(1\right)}}\end{array}\right)\,\tilde{\psi}_{22}^{T}=\left(\begin{array}[]{c}{\tilde{e}_{W}^{\left(2\right)}}\\ {\tilde{\nu}_{W}^{\left(2\right)}}\\ {\tilde{\nu}^{\left(2\right)}}\end{array}\right)\,\tilde{\psi}_{23}^{T}=\left(\begin{array}[]{c}{\tilde{e}_{W}^{\left(3\right)}}\\ {\tilde{\nu}_{W}^{\left(3\right)}}\\ {\tilde{\nu}^{\left(3\right)}}\end{array}\right).$
In the above labelling, lower numerical indices are fundamental-representation indices for the unbroken SU(3) group (the strong interaction). Fermions with a “W” index interact with the W boson (as members of an isodoublet). Based on the magnitude of their electric charges, up-type quarks, down-type quarks, charged leptons, and neutral leptons are labelled with $u,d,e,\nu$. All of the fermion fields are 2-component Weyl fermions with a lower, undotted spin index. In the convention of [7] (which is also the Wess/Bagger & Bilal convention of [15]), a Weyl fermion with a lower undotted index corresponds to a right-chiral fermion that vanishes when acted on by $1-\gamma_{5}$ (see appendix A of [16]).
In that convention, the $u$ and $d$ fields are right-chiral fermions with electric charges of ${\tfrac{2}{3}}$ and $-{\tfrac{1}{3}}$, respectively, so they are mapped to right-chiral quarks. The fields $\tilde{u}$ and $\tilde{d}$ are right-chiral fermions with electric charges of $-{\tfrac{2}{3}}$ and ${\tfrac{1}{3}}$, respectively, so they are mapped to Hermitian conjugates of left-chiral quarks. One way that this model differs from the Standard Model is that some of the right-chiral quarks have a “W” index so they interact with the W boson, while some of the left-chiral quarks lack that index so they do not interact with the W boson. That difference from the Standard Model will be discussed further in the next section of this paper.
Mapping of leptons in this model is described in the next section, but one case is mentioned here: The right-chiral fields $e_{W}^{\left(1\right)}$ and $\nu_{W}^{\left(1\right)}$ (with electric charges of $+1$ and 00) are mapped to $e_{L}^{\dagger}$ and $\nu_{e}^{\dagger}$, Hermitian conjugates of the left-chiral electron and its neutrino. The right-chiral field $\tilde{e}^{\left(1\right)}$ (with charge $-1$ and no $W$ subscript) is mapped to $e_{R}$, the right-chiral electron. Since only fields with a $W$ subscript connect to the W boson, this convention ensures that only the left-chiral electrons (and their Hermitian conjugates) interact with the W boson.
3 Masses and Mixing of Observed Particles
This section begins by discussing quantum-generated interactions and their effect on the vacuum. Given certain assumptions about those quantum interactions, it is shown how this model produces the observed spectrum of particle masses. To validate the assumptions made, detailed quantum calculations would be needed, and those calculations are not performed in this paper. Instead, a picture is sketched as to what those calculations would need to produce in order to generate measured masses and mixing.
3.1 Confinement
For a supersymmetric gauge theory involving a chiral superfield in the adjoint representation, it has been shown that a tree-level mass term $m_{\Phi}$ for the adjoint superfield will lead to quark confinement [17, 18, 15, 14]. The duality inherent in these theories permits moving from a description in terms of strongly coupled scalars with color-electric charge to a description in terms of weakly coupled monopoles with color-magnetic charge. A tree-level $m_{\Phi}$ mass term can cause the vacuum to settle on one of two configurations where the vev of the trace of the square of the adjoint superfield does not vanish $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle\neq 0$. In one of those configurations, color-magnetic monopoles become massless, condense, and cause quark confinement through a dual Meisner effect [17, 18, 15, 14].
The theory of this paper is not supersymmetric, since its “superfields” are “twisted”. Nonetheless, this theory does have an adjoint-representation “twisted superfield” with a mass $m_{\Phi}$. Also, the scalars in this theory are in the same representation as the scalars in the corresponding supersymmetric theory, so the vacuum moduli spaces of the two theories should be similar, particularly below the scale where the $m=1$ fundamental and conjugate scalars of the theory get large masses (see section 5). That being the case, it is speculated that quantum effects similar to those in the supersymmetric theory cause the following vev to form:
$\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle\neq 0\,\,\,\textrm{quantum mechanically}.$ (3.1)
It is further speculated that the similarity with the corresponding supersymmetric moduli space is sufficient so that the scalars in $\varphi_{1}$ form color-magnetic monopoles that become massless, condense and cause quark confinement through the dual Meisner effect.
The vev $\left\langle\varphi_{1}\right\rangle$ will also generate quark mass terms of the following form (connecting gaugino and adjoint quarks):
$2\sqrt{2}g_{3}i\int d^{4}x\rm{tr}\left(\tilde{\chi}\left\langle\varphi_{1}^{\dagger}\right\rangle\lambda-\tilde{\lambda}\left\langle\varphi_{1}^{\dagger}\right\rangle\chi\right)+h.c.,$ (3.2)
where it has been assumed that the strong coupling $g_{3}$ is the appropriate coupling to use for these quark mass terms.
Before concluding this subsection, there are a couple of additional quantum effects to be discussed. For supersymmetric theories, instantons can generate nonperturbative low-energy effective superpotential terms. The appendix proposes that a similar effect occurs for this theory. To accommodate these terms, it is assumed that the following vevs that are zero classically acquire small but nonzero vevs quantum mechanically:
$\bar{\tilde{\phi}}_{21},\bar{\phi}_{22}\neq 0\textrm{ quantum mechanically}.$ (3.3)
These vevs set the scale for neutrino masses later in this section.
As discussed in the appendix, the quantum-generated superpotential terms also try to make $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$ small. To accomplish this in the presence of eq (3.1), it is assumed that the quantum vacuum modifies $\left\langle\varphi_{2}\right\rangle$ of eq (2.6) to the following:
$\left\langle\varphi_{2}\right\rangle=\frac{i\bar{\varphi}_{2}}{\sqrt{2\left(1+c^{2}\right)}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{c}\\ {0}&{1}&{0}\end{array}\right),$ (3.4)
where the quantum-generated $c$ is assumed to be much smaller than 1. The square of the above vev has a nonvanishing trace that can reduce $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$.
3.2 Observed boson masses
From the symmetry breaking defined in eq (2.9), the mass of the $W$ boson is primarily determined by $\bar{\phi}_{21}^{2}$. In other words, at tree level:
$M_{W}^{2}\simeq{\tfrac{1}{2}}g_{2}^{2}\bar{\phi}_{21}^{2}.$ (3.5)
Phenomenologically, the mass of the W boson determines the vev $\bar{\phi}_{21}^{2}$, with the running coupling $g_{2}^{2}$ evaluated at the W boson mass scale.
The Z boson mass in this model is also primarily determined by $\bar{\phi}_{21}^{2}$. Due to the form of eq (2.16), the Z boson mass in this model differs slightly from the Standard Model expression. Namely, the mass of the Z boson is:
[…]
where the angles are defined by coupling constants as in eq (2.17). In section 5 when discussing the $Z$ boson width, a scenario is presented where $\sin\phi_{Z}\simeq 0.1$ at the $Z$ boson scale (getting much smaller at the $Z^{\prime}$ boson scale). The implications for precision $W$ mass measurements are discussed in that section. In most scenarios, $\sin\phi_{Z}$ should not be larger than that (and possibly much smaller), so the following approximation can be used for most purposes in this paper:
$\cos\theta_{Z}$ $\simeq\cos\theta_{W}$ (3.7)
$\cos\phi_{Z}$ $\simeq 1,$
where $\theta_{W}$ is the Weinberg angle.
In this model, the scalar vev $\bar{\phi}_{21}$ also generates most of the mass of the observed Brout-Englert-Higgs boson, through the d-term part of the scalar potential. The d-term has two parts: one part from terms like $\left\langle d_{2}^{3}\right\rangle d_{2}^{3}$, and the other part where each $d_{2}^{A}$ in $\tfrac{1}{2}d_{2}^{A}d_{2}^{A}$ has one vev and one Higgs field. Those d-term contributions can be found by first re-expanding the d term part of the scalar potential of eq (2.1) using the following basis of diagonal U(3)$\times$×U(3) generators:
$g_{U}T^{U},g_{Y}T^{Y},g_{Y^{\prime}}T^{Y^{\prime}},g_{2}T_{2}^{3},g_{3}T_{1}^{8},g_{3}T_{1}^{3}.$ (3.8)
The expansion of the complex scalar field $\phi_{21}$ into its component fields includes the following:
$i\phi_{21}=\left(\bar{\phi}_{21}+{\tfrac{1}{\sqrt{2}}}h_{21}\right)\left(1,0,0\right)^{T}+...,$ (3.9)
where $h_{21}$ is the real scalar Higgs boson field and ++...+ … includes the fields that get “eaten” by the W and Z gauge bosons.
In the basis of eq (3.8), the vevs of the auxiliary d fields at the minimum of the scalar potential are approximately:
$-g_{U}\left\langle d_{U}\right\rangle$ $\simeq\sqrt{2}m_{21}^{2}$ (3.10)
$-g_{2}\left\langle d_{2}^{3}\right\rangle$ $=\tfrac{1}{2}g_{2}^{2}\bar{\phi}^{2}_{21}$
$-g_{Y}\left\langle d_{Y}\right\rangle$ $=\tfrac{3\sqrt{3}}{2\sqrt{5}}g_{Y}^{2}\bar{\phi}^{2}_{21}\,,$
where it has been assumed that $g_{+}=g$ at the unification scale and that $g_{Y^{\prime}}$ is negligible due to reasons discussed below. The approximation has also assumed that $\bar{\phi}^{2}_{21}\gg\bar{\tilde{\phi}}_{22},\bar{\varphi}_{2}$.
Assuming $m_{211}$ is small, the mass of the Higgs boson (before radiative corrections) is:
$M_{H}^{2}$ $\simeq m_{21}^{2}-\tfrac{1}{2}g_{2}\left\langle d_{2}^{3}\right\rangle+\tfrac{\sqrt{3}}{2\sqrt{5}}g_{Y}\left\langle d_{Y}\right\rangle+\tfrac{1}{3\sqrt{2}}g_{U}\left\langle d_{U}\right\rangle$ (3.11)
$+2\bar{\phi}_{21}^{2}\left(g_{U}^{2}\left(T_{1}^{U}\right)_{44}^{2}+g_{Y}^{2}\left(T_{1}^{Y}\right)_{44}^{2}+g_{2}^{2}\left(T_{2}^{3}\right)_{44}^{2}\right)$
$\simeq M_{Z}^{2}+(\tfrac{11}{18}-\tfrac{3}{2}\tan^{2}\theta_{W})M_{W}^{2}+\tfrac{2}{3}m_{21}^{2}$
$\simeq\left(97\,{\rm GeV}\right)^{2}+\tfrac{2}{3}m_{21}^{2},$
where the scalar mass $m_{21}$ is from eq (1.18). The fourth line above comes from the relations in eqs (2.17) and (3.6) along with the approximation that $g_{U}\simeq g_{2}$. In the $Z$ width part of section 5, it is suggested that $m_{21}$ may be very small. In that case, roughly 28 GeV of the Higgs Boson mass in this model is assumed to come from radiative corrections.
The $h_{21}$ real scalar field described above has the same interactions with the W boson, Z boson, and top quark as does the Standard Model Higgs boson. So this model is consistent with measurements of Higgs boson decays and interactions involving these particles, since those measurements are consistent with the Standard Model.
On the other hand, interactions of $h_{21}$ with other quarks or leptons are different in this model than in the Standard Model. As described later in this section, leptons and quarks (other than the top) acquire masses from different mechanisms, not from Yukawa interactions with the Higgs Boson. Those particles do have interactions with $h_{21}$ via the nonperturbative terms described in the Appendix. More work would need to be done to determine whether those interactions were consistent with observations.
3.3 Quark masses and mixing
Keeping in mind the particle designations of this model defined in eq (2.21), the up-type quarks can be arranged into the following 3x3 mass matrix (and its Hermitian conjugate):
$\rm{Rows:}\,\tilde{u}^{\left(1\right)},\tilde{u}_{W}^{A},\tilde{u}_{W}^{G}\,\,\,\rm{Columns:}\,u_{W}^{G},u_{W}^{A},u^{\left(1\right)}$ (3.12)
$\mathit{M}_{u}=\left(\begin{array}[]{ccc}{\hat{g}\bar{\tilde{\phi}}_{21}}&{\tfrac{1}{\sqrt{2}}\tilde{\Gamma}_{211}\bar{\phi}_{21}}&{\tilde{m}_{211}}\\ {\tilde{M}_{G}^{(1)}-\Delta}&{\tilde{m}_{\Phi}}&{\tfrac{1}{\sqrt{2}}\bar{\tilde{\phi}}_{21}\tilde{\Gamma}_{211}}\\ {\tilde{m}_{\lambda}}&{\tilde{M}_{G}^{(1)}+\Delta}&{\hat{g}\bar{\phi}_{21}}\end{array}\right),$
where the gaugino coupling $\hat{g}$ is discussed below in eq (3.14). A tilde is put on tree-level masses and superpotential couplings to show that they include quantum modifications from the appendix. For all of the non-neutrino fermion mass matrices in this paper (including the above matrix) the rows have an upper undotted spin index (using the convention of [7]) while the columns have a lower undotted spin index (e.g. $\tilde{u}^{\left(1\right)\alpha}u_{W\alpha}^{G}$). Those indices are summed over, and they are suppressed.
The parameter $\tilde{M}_{G}^{(1)}$ is generated nonperturbatively from terms like eq (A.12) in the Appendix. The mass $\Delta$ is from eq (3.2).
The first, second, and third generation up-type quarks correspond to the first, second and third rows and columns of the up-type quark matrix. To a first approximation, the third-generation quark is just the top quark and its mass is approximately:
$m_{t}\simeq\hat{g}\bar{\phi}_{21}\,\,\textrm{Top quark}.$ (3.13)
Recalling the convention from eq (2.21) that fields with a tilde are Hermitian conjugates of left-chiral quarks while those without one are right-chiral quarks, it can be seen from eq (3.12) that for the third-generation quark, only its left-chiral component interacts with the W boson. That allows this model to be consistent with top-quark polarization measurements by ATLAS [19].
The first- and second-generation quark interactions with the W boson, however, differ from those of the Standard Model. For the second-generation up-type quark (mostly charm), both its left- and right-chiral components interact with the W boson. For the first-generation up-type quark (mostly up), only its right-chiral component interacts with the W boson. Nuclear beta decay experiments are consistent with the W boson connecting left-chiral protons to left-chiral neutrons. If the spin of a proton was primarily just the sum of the spins of its three valence quarks, then the right-chiral W boson interactions of this model’s up (and down) quarks might not be consistent with nuclear beta decay experiments. However, the spin of the proton mostly comes from gluons and orbital angular momentum [20], so this model’s right-chiral interactions are not inconsistent with nuclear beta decay data.
Above in eq (3.5), it was shown that the magnitude of the vev $\bar{\phi}_{21}$ is approximately determined by the W boson mass and the SU(2) weak coupling constant $g_{2}$. In this model, the top quark mass is also determined by $\bar{\phi}_{21}$, but multiplied by the gaugino coupling $\hat{g}$ rather than the weak coupling $g_{2}$ (at the unification scale, these are the same). To be consistent with observation, this model would need to show that the difference in the coupling constants $g_{2}$ vs. $\hat{g}$ at the electroweak scale is the amount needed for the model to correctly reproduce both the W boson mass and the top quark mass.
To get the actual value for this model’s top quark mass, the calculation should be performed to determine how $\hat{g}$ runs in this model as the scale is lowered from the unification scale. Such a calculation is outside the scope of this paper.
Instead, the following approach is used: Fits are made below to quark masses and CKM data. Those fits produce a numerical value for $\hat{g}\bar{\phi}_{21}$. By comparing that numerical value to the mass of the W boson given by $\sqrt{2}M_{W}\simeq g_{2}\bar{\phi}_{21}$, the following ratio can be extracted:
$\hat{g}\simeq 1.5g_{2}\simeq 0.8g_{3},$ (3.14)
where $g_{2}$ and $g_{3}$ are evaluated at the Z boson mass scale. It is presumed that the gaugino coupling $\hat{g}$ of this model takes the above value at the Z boson mass scale.
An interesting comparison is provided by one of the results from the Minimal Supersymmetric Standard Model [21, 22, 23]. In the MSSM, the Yukawa coupling for the top quark can be approximated by [11]
$\lambda_{t}^{2}\left(M_{Z}\right)$ $=\frac{7g_{3}^{2}\left(M_{Z}\right)}{18}\left(1-\left(\frac{\alpha_{3}\left(M_{U}\right)}{\alpha_{3}\left(M_{Z}\right)}\right)^{{7\mathord{\left/{\vphantom{79}}\right.\kern-1.2pt}9}}\right)^{-1}$ (3.15)
$\sim\left(0.7g_{3}\right)^{2},$
where $M_{U}$ is the unification scale (and 7/18 is called the Pendleton Ross fixed point). The value of the strong coupling at the unification scale in the above formula is taken from section 4 of this paper.
Before wrapping up this discussion of up-type quarks, it is worth repeating that in this model, the one vev $\bar{\phi}_{21}$ chosen by the vacuum is presumed to determine not only the masses of the W, Z and Higgs bosons, but also the mass of the top quark without the addition of a separate Yukawa coupling.
From eq (2.21), it can be seen that there are twice as many flavors of down-type quarks in this model $(6L\times 6R)$ as there are up-type quarks $(3L\times 3R)$. However, the unification scale vevs generate unification-scale masses for $\tilde{d}^{G}d^{\left(3\right)}$ and $\tilde{d}^{\left(3\right)}d^{G}$, so those down-type quarks decouple. The remaining 4x4 down-type quark mass matrix has the following structure:
$\rm{Rows:}\,\tilde{d}^{A},\tilde{d}_{W}^{A},\tilde{d}_{W}^{G},\tilde{d}^{\left(2\right)}\,\,\,\rm{Columns:}\,d_{W}^{G},d^{\left(2\right)},d^{A},d_{W}^{A}$ (3.16)
$M_{d}=\left(\begin{array}[]{cccc}{...}&{...}&{\tilde{m}^{\prime\prime}_{\Phi}}&{\hat{g}\bar{\varphi}_{2}}\\ {...}&{\tfrac{1}{\sqrt{2}}\bar{\tilde{\phi}}_{22}\tilde{\Gamma}_{222}}&{\tfrac{1}{\sqrt{2}}\bar{\varphi}_{2}\tilde{\Gamma}_{\Phi}}&{\tilde{m}^{\prime}_{\Phi}}\\ {\tilde{m}^{\prime}_{\lambda}}&{...}&{...}&{...}\\ {\hat{g}\bar{\tilde{\phi}}_{22}}&{\tilde{m}_{222}}&{...}&{...}\end{array}\right),$
where only tree-level couplings are shown explicitly. Tildes again indicate that these couplings can be modified by quantum-generated couplings discussed in the appendix. Quantum couplings can also be generated for the components without tree-level contributions; those are denoted by “...”. Primes and double primes on $\tilde{m}^{\prime}_{\Phi}$ and $\tilde{m}^{\prime}_{\lambda}$ are used to show that these masses in the down-type matrix do not need to be the same as those in the up-type matrix, as discussed in the Appendix.
Construction of the CKM matrix in this model is more involved than in the Standard Model. As in the Standard Model, the first step is to find unitary matrices $V_{L}^{U}$, $V_{R}^{U}$, $V_{L}^{D}$, and $V_{R}^{D}$ that diagonalize the quark mass matrices:
$V_{L}^{U}\mathit{M}_{u}V_{R}^{U{\dagger}}\,\,\rm{and}$ (3.17)
$V_{L}^{D}\mathit{M}_{d}V_{R}^{D{\dagger}}\,\,\textrm{are diagonal.}$.
Unlike in the Standard Model where only the $V_{L}^{U,D}$ matrices have connections to the W boson, in this model, all four of the matrices have connections to the W.
From the diagonalizing matrices, one may construct the following 2 versions of CKM matrices:
$V_{\rm{CKM}}^{\pm}$ $=V_{R}^{U}\left(\begin{array}[]{cccc}{1}&{0}&{0}&{0}\\ {0}&{0}&{0}&{1}\\ {0}&{0}&{0}&{0}\end{array}\right)V_{R}^{D{\dagger}}$ (3.18)
±$\pm V_{L}^{U}\left(\begin{array}[]{cccc}{0}&{0}&{0}&{0}\\ {0}&{1}&{0}&{0}\\ {0}&{0}&{1}&{0}\end{array}\right)V_{L}^{D{\dagger}}$
The placement of the 1’s in the above matrices is based on which quark fields in eqs (3.12) and (3.16) have a $W$ subscript (signifying that they interact with the W boson to change an up-type quark to a down-type quark and vice versa). In this model the “CKM” matrix is not a square, it is 3x4.
Also, in this model, a different CKM matrix should be used depending on whether a vector current or axial vector current process is being considered. Specifically:
Vector current decays: $V_{\rm{CKM}}^{+}$ (3.19)
Axial vector current decays: $V_{\rm{CKM}}^{-}.$
For CKM measurements involving vector current decays (e.g. an exclusive semi-leptonic decay from one spin-0 meson to a different spin-0 meson), $V_{\rm{CKM}}^{+}$ should be used for comparison to this model. For CKM measurements involving axial vector current decays (e.g. the purely leptonic decay of a spin-0 pseudo-scalar meson), $V_{\rm{CKM}}^{-}$ should be used.
To fit the parameters in the up- and down-type quark mass matrices, it is first assumed that the fourth down-type quark has a mass of 2.8 GeV. The reasons and experimental justifications for this choice are discussed in section 5.
A fit of this model’s parameters was performed using the six known quark masses, this new quark mass, and measured CKM data, both for vector-current and axial-vector-current decays. The following values (in GeV) were obtained for elements of the mass matrices:
$\rm{Rows:}\,\tilde{u}^{\left(1\right)},\tilde{u}_{W}^{A},\tilde{u}_{W}^{G}\,\,\,\rm{Columns:}\,u_{W}^{G},u_{W}^{A},u^{\left(1\right)}$ (3.20)
$\mathit{M}_{u}=\left(\begin{array}[]{ccc}{0.00131}&{-0.00191}&{0.0122}\\ {0.6493}&{1.1126}&{-5.3477}\\ {1.7379}&{-1.0843}&{171.94}\end{array}\right).$
$\rm{Rows:}\,\tilde{d}^{A},\tilde{d}_{W}^{A},\tilde{d}_{W}^{G},\tilde{d}^{\left(2\right)}\,\,\,\rm{Columns:}\,d_{W}^{G},d^{\left(2\right)},d^{A},d_{W}^{A}$ (3.21)
$\mathit{M}_{d}=\left(\begin{array}[]{cccc}{-0.0338}&{0.0137}&{0.0246}&{-0.0527}\\ {-0.0252}&{-0.0725}&{-0.0636}&{-0.0804}\\ {0.1054}&{1.4894}&{-3.8554}&{0.1329}\\ {1.4717}&{-0.0800}&{0.2408}&{2.3052}\end{array}\right).$
To find the unitary matrices that diagonalized the above matrices and generated CKM matrices, each mass matrix was first multipled by its transpose, then matrices of left- and right- eigenvectors were found. The allowed flexibility to introduce complex phases was not used for the fits of this paper, so no attempt was made to fit the experimentally measured complex phases of the CKM matrix.
With the above mass matrices, the model generates the following for the quark masses and the CKM matrices:
$m_{u}=0.002,\,\,\,m_{c}=1.28,\,\,\,m_{t}=172$ (3.22)
$m_{d}=0.005,\,\,\,m_{s}=0.095,\,\,\,m_{b}=4.18,\,\,\,m_{f}=2.8$
$\left|V_{\rm{CKM}}^{+}\right|=\left(\begin{array}[]{cccc}{0.9745}&{0.2239}&{0.0043}&{0.0094}\\ {0.2323}&{0.9661}&{0.0388}&{0.9720}\\ {0.0088}&{0.0402}&{0.9973}&{0.0619}\end{array}\right)$
$\left|V_{\rm{CKM}}^{-}\right|=\left(\begin{array}[]{cccc}{0.9742}&{0.2252}&{0.0042}&{0.0093}\\ {0.2165}&{0.9805}&{0.0394}&{1.0278}\\ {0.0144}&{0.0349}&{0.9974}&{0.0621}\end{array}\right),$
where all mass values are in GeV, and the subscript $f$ was used to denote the fourth down-type quark.
[…]
3.4 Charged leptons
After unification scale symmetry breaking but before electroweak symmetry breaking, this model has the following lepton content: (i) three fundamental-representation isodoublet pairs, (ii) three conjugate-representation isodoublet pairs, (iii) a fundamental and a conjugate isosinglet charged lepton, and (iv) 2 fundamental and 2 conjugate isosinglet neutrinos. Isosinglet leptons do not interact with the W boson.
This section maps this model’s charged leptons to the observed leptons. The first part of this section maps the flavor 1 charged leptons to the electron and an as-yet-undiscovered heavy lepton. The second part maps the flavor 2 and 3 charged leptons to the muon and tau lepton.
From eqs (A.10), (A.16) and (A.17) of the appendix, the flavor-1 lepton mass matrix (for Dirac masses) takes the following form:
$\begin{array}[]{c}{\left(\begin{array}[]{ccc}{\tilde{\nu}_{W}^{\left(1\right)}}&{\tilde{e}_{W}^{\left(1\right)}}&{\tilde{e}^{\left(1\right)}}\end{array}\right)}\\ {}\hfil\\ {}\hfil\end{array}\left(\begin{array}[]{ccc}{\tilde{m}_{111}}&{0}&{0}\\ {0}&{\tilde{m}^{\prime}_{111}}&{M_{\Omega}}\\ {0}&{m_{e}}&{\tilde{m}^{\prime}_{111}}\end{array}\right)\left(\begin{array}[]{c}{\nu_{W}^{\left(1\right)}}\\ {e_{W}^{\left(1\right)}}\\ {e^{\left(1\right)}}\end{array}\right)\,,$ (3.23)
where $m_{e}$ is the mass of the electron and $M_{\Omega}$ is the mass of a new, proposed heavy lepton. This mass should be in the range 103 GeV $<M_{\Omega}<2\times 10^{6}$ GeV. The lower limit comes from current 95% confidence exclusion limit from heavy lepton searches [26]. The upper limit is from the anomaly-based analysis surrounding eq (4.1).
From eq (2.19), the last three diagonal components of the photon field when acting on flavor-1 fundamental-representation fields are $e\left(0,1,1\right)$. As a result, if the left- and right-chiral electron components are defined via $e_{L}^{\dagger}=e_{W}^{\left(1\right)}$ and $e_{R}=\tilde{e}^{\left(1\right)}$, then those fields have the correct electric charges of $+1$ and $-1$ as well as the correct isodoublet and isosinglet designations. In the limit that $\tilde{m}^{\prime}_{111}\ll m_{\Omega}$, the flavor 1 leptons can therefore be identified as follows:
$\begin{array}[]{ccc}{\nu_{W}^{(1)}=\nu_{eL}^{\dagger}}&{e_{W}^{(1)}=e_{L}^{\dagger}}&{e^{(1)}=\Omega_{L}^{\dagger}}\\ {\tilde{\nu}_{W}^{(1)}=\nu_{\Omega R}}&{\tilde{e}_{W}^{(1)}=\Omega_{R}}&{\tilde{e}^{(1)}=e_{R},}\end{array}$ (3.24)
where $\Omega$ is the hypothesized new heavy lepton. From this identification, it can be seen that the right-chiral component (not left-chiral) of the negatively charged Omega lepton is the one that interacts with the W boson, forming an isodoublet with its right-chiral neutrino.
The electron not only has the same interactions with the photon and W boson as in the Standard Model, it also has the same interactions with the Z boson, as can be seen from eqs (2.18) and (2.20).
In this model, the muon and tau lepton come from the flavor 2/3 lepton sector of the theory. From eqs (A.10), (A.16) and (A.17) of the appendix, the flavor 2/3 lepton mass matrix takes the form:
$\left(\begin{array}[]{c}{\tilde{e}_{W}^{\left(F\right)}}\\ {\tilde{\nu}_{W}^{\left(F\right)}}\\ {\tilde{\nu}^{\left(F\right)}}\end{array}\right)^{T}\left(\begin{array}[]{ccc}{\tilde{m}_{1FF^{\prime}}}&{0}&{0}\\ {0}&{\tilde{m}^{\prime}_{1FF^{\prime}}}&{\tilde{m}_{\nu FF^{\prime}}}\\ {0}&{m_{\nu FF^{\prime}}}&{\tilde{m}^{\prime}_{1FF^{\prime}}}\end{array}\right)\left(\begin{array}[]{c}{e_{W}^{(F^{\prime})}}\\ {\nu_{W}^{(F^{\prime})}}\\ {\nu^{(F^{\prime})}}\end{array}\right),$ (3.25)
with $F,F^{\prime}=2,3$. As will be discussed below, the right-chiral $\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$ are mapped to “heavy” right-chiral neutrinos with masses similar to that of the tau lepton. The right-chiral $\tilde{\nu}_{W}^{\left(F\right)}$ are mapped to two of the the standard antineutrinos.
Consistent with those mappings, the right-chiral leptons $e_{W}^{(F)}$ are mapped to some combination of $\mu_{R}$ and $\tau_{R}$, since they behave like isosinglets for $W$-mediated decays of particles with mass less than $m_{\tau}$. The right-chiral leptons $\tilde{e}_{W}^{\left(F\right)}$ are mapped to some combination of $\mu_{L}^{\dagger}$ and $\tau_{L}^{\dagger}$, making them isodoublets with the light antineutrinos. These mappings also impart the correct electric charge since from eq (2.19), the last three components of the photon for flavor 2,3 fields are $e\left(-1,0,0\right)$.
The Z-boson interactions of the muon and tau lepton of this model are different from those of the Standard Model; they have only vector (not axial vector) interactions. It will be shown in section 5 that this feature of the model leads to a different value for the anomalous magnetic moment of the muon than generated by Standard-Model calculations. On the other hand, the feature makes it challenging to reproduce the precision $e^{+}e^{-}$ data from LEP and SLD that were consistent with the muon and tau lepton having the Standard-Model coupling to the Z boson. It will be shown in section 5 how the “heavy” right-chiral neutrinos mentioned above provide a way for the model to reproduce LEP and SLD data.
Since the mass parameters $\tilde{m}_{1FF^{\prime}}$ of eq (3.25) are not diagonal in 2/3 flavor space, the flavor space of this paper does not coincide with the muon and tau lepton mass eigenstates. In 2/3 flavor space, the mass matrix for the charged leptons can be diagonalized as follows:
$\left(\begin{array}[]{cc}{\tilde{m}_{122}}&{\tilde{m}_{123}}\\ {\tilde{m}_{132}}&{\tilde{m}_{133}}\end{array}\right)=\tilde{u}_{23}^{{\dagger}}\left(\begin{array}[]{cc}{m_{\mu}}&{0}\\ {0}&{m_{\tau}}\end{array}\right)u_{23}$ (3.26)
$u_{23}=\left(\begin{array}[]{cc}{\cos\omega_{23}}&{\sin\omega_{23}}\\ {-\sin\omega_{23}}&{\cos\omega_{23}}\end{array}\right)$
$\tilde{u}_{23}=\left(\begin{array}[]{cc}{\cos\tilde{\omega}_{23}}&{\sin\tilde{\omega}_{23}}\\ {-\sin\tilde{\omega}_{23}}&{\cos\tilde{\omega}_{23}}\end{array}\right),$
where $m_{\mu}$ and $m_{\tau}$ are the masses of the muon and tau lepton.
The angles $\omega_{23}$ and $\tilde{\omega}_{23}$ define the rotations from the charged lepton mass eigenbasis to the 2/3 flavor basis of this paper. In particular, $\tilde{\omega}_{23}$ defines the rotation that the light flavor 2/3 neutrinos must undergo to transform from the muon/tau mass basis to the flavor basis. It is worth noting that the 2/3 flavor basis cannot just be redefined to coincide with the charged lepton mass eigenbasis, since the unification scale vevs discussed in the last section have already defined the flavor “3” direction.
3.5 Neutrinos
It has been well established experimentally that the 3 known light neutrinos have very small (sub-$eV$) masses, and that their mass eigenbasis is different from the mass eigenbasis of the charged leptons (electron, muon, tau). The 3$\times$3 Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix describes the rotation between these two mass eigenbases. The three angles and one phase describing this unitary matrix have been determined experimentally (see for example [27]).
This model’s neutrino mass eigenbasis is also different from its charged lepton mass eigenbasis. Multiple rotations generate this difference: the above-described flavor “23” rotation from the charged lepton mass basis to the flavor basis of this paper and additional rotations needed to diagonalize the neutrino mass matrix from the form it takes in the flavor basis. By phenomenologically choosing a combination of tree-level and quantum-generated parameters, there are many ways for the model to reproduce both the neutrino mass differences and mixing angles observed experimentally.
From eqs (3.23) and (3.25), the Dirac matrix connecting the 10 neutrinos of the model can be very complicated. To illustrate mechanisms that can give rise to observed neutrino masses and mixing, the following simplified Dirac matrix will be considered instead:
[…]
The rows of the matrix are right-chiral neutrinos and the columns are Hermitian conjugates of left-chiral neutrinos. The first three columns are the observed 3 light antineutrinos in the flavor basis of this paper.
In eq (A.11) of the appendix, it is argued that nonperturbative quantum effects generate a large Majorana mass $\tilde{M}_{\nu}$ for the following linear combination of right-chiral neutrinos:
$\tilde{\nu}_{M}=\cos\theta_{12}\tilde{\nu}_{W}^{(1)}+\sin\theta_{12}\nu_{W}^{(2)},$ (3.28)
where $\tan\theta_{12}=\bar{\tilde{\phi}}_{21}/\bar{\phi}_{22}$. The first two rows of eq (3.27) can be rotated to the basis of $\tilde{\nu}_{M}$ and its orthogonal combination $\nu_{M}$, with $\tilde{\nu}_{M}$ in the first row. For $\sin\theta_{12}\ll 1$, $\tilde{m}^{\prime}_{122}\gg\tilde{m}^{\prime}_{111}$ and $\tilde{m}^{\prime}_{122}\gg m_{\nu 2}$, a significant rotation between the first and second columns together with a very small rotation between the first and fifth columns could be necessary to again make all of the entries of the first row vanish except the “11” element.
This simple example has shown how at least 2 rotations may be necessary to transform to the neutrino mass eigenbasis: a “23” rotation to get from the charged lepton mass basis to the flavor basis and a “12” rotation to get from the flavor basis to the heavy Majorana neutrino basis. Even in this simple model, additional rotations will be necessary for full diagonalization. The parameters of the model may be chosen so that the net result of all of the rotations is the PMNS matrix.
The mass eigenvalues of the matrix can have very disparate values. For example, if the equality $m^{2}_{133}=m_{\nu 3}\tilde{m}_{\nu 3}$ is assumed with $m_{\nu 3}\gg\tilde{m}_{\nu 3}$, then the flavor-3 part of the neutrino mass matrix of eq (3.27) has zero determinant, involving a large eigenvalue and a zero eigenvalue. A slight deviation from that mass equality can generate a mass for $\sim\nu_{W}^{(3)}$ that is near the tau lepton mass and a mass for $\sim\tilde{\nu}_{W}^{(3)}$ that is around 0.05 eV. A similar mechanism for the flavor-2 neutrinos could result in neutrino masses near $m_{\tau}$ and 0.0086 eV. Finally a seesaw mechanism involving the large Majorana mass of $\tilde{\nu}_{M}$ could generate a neutrino mass much smaller than 0.0086 eV. Those mechanisms would then reproduce the observed neutrino mass differences after PMNS diagonalization from the charged lepton mass basis to the light neutrino mass basis.
As mentioned above, this model predicts two right-chiral neutrinos that are isodoublets with the right-chiral muon and tau lepton. Both of these neutrinos are assumed to have masses near that of the tau lepton. The implications of these neutrinos are discussed in section 5.
4 Unification and Anomalies
As mentioned in section 1, at a scale above any symmetry breaking, the model of this paper is free of anomalies. The reason is simple: for every fermion in every representation, there is another fermion in a conjugate representation of each of the SU(3) groups that also has opposite charges for the Abelian fields $A_{+\mu}^{0}$ and $A_{-\mu}^{0}$. Consequently, all of the gauge anomaly triangle diagrams cancel.
The unification scale $M_{U}$ of this model is defined as the scale where the vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ break the original SU(3)$\times$×SU(3)$\times$×U(1)$\times$×U(1) symmetry down to SU(3)$\times$×SU(2)$\times$×U(1)$\times$×U(1). Below the unification scale, the four couplings $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$ for these four groups run differently. As mentioned in section 3, this symmetry breaking causes the quark pairs $\tilde{d}^{G}d^{\left(3\right)}$ and $\tilde{d}^{\left(3\right)}d^{G}$ to acquire unification-scale masses. These heavy fermions are in conjugate representations of the remaining SU(3) group, are SU(2) singlets, and have equal and opposite charges for the Abelian $Y$ and $Y^{\prime}$ fields. Consequently, the effective theory below the unification scale that ignores these heavy quarks is also free of local gauge anomalies.
As can be seen from eq (2.15), gauginos below the unification scale have $Y^{\prime}$ charges that are $2\sqrt{6}$ times larger than their $Y$ charges. From eq (1.20), adjoint-representation quarks have an even larger ratio of their $Y^{\prime}$ charges to their $Y$ charges. The $Y^{\prime}$ charges of all of the other fermions and scalars are also larger than their $Y$ charges. The net result is that below the unification scale, $g_{Y^{\prime}}$ drops to lower scales much more quickly than $g_{Y}$. This is a partial justification for the small value of $\sin\phi_{Z}$ assumed in sections 3 and 5.
Below the unification scale, the next lower scale of the theory is defined by the large Majorana mass $\tilde{M}_{\nu}$ of the right-chiral neutrino $\tilde{\nu}_{M}$ discussed in eq (A.11) of the Appendix. An effective theory below the scale of $\tilde{M}_{\nu}$ that excluded that neutrino would have an anomaly (see [28, 29]). Applying arguments from [30], this effective anomaly implies the following upper limit for the mass of that neutrino:
$\tilde{M}_{\nu}\lesssim\frac{32\pi^{3}M_{Z}}{|\tfrac{1}{2}\sqrt{\tfrac{3}{5}}g_{Y}|^{3}}\simeq 2\times 10^{6}\textrm{ GeV}.$ (4.1)
When $\tilde{M}_{\nu}$ obeys the above relation, the $Z$ boson mass is able to resolve the obstruction to renormalizability of the effective theory that is generated by anomalous diagrams involving the $Y$ boson.
As discussed in section 2, after the $Z$ and $W$ bosons acquire masses at the electroweak scale, the remaining massless bosons are gluons, photons, and the $Z^{\prime}$ boson. Gluons and photons do not couple to the $\tilde{\nu}_{M}$ heavy neutrino, so they do not generate an anomaly. However, the $Z^{\prime}$ boson does couple to it. Since the $Z^{\prime}$ boson does not acquire mass at the electroweak scale, the only way for it to satisfy a relation analogous to that of eq (4.1) is by $g_{Y^{\prime}}$ (and consequently $g_{Z^{\prime}}$) being driven to a very small value.
Said another way, since the underlying theory with the heavy neutrino is anomaly free, the $Z^{\prime}$-induced anomaly of the effective theory must manifest itself by causing the $Z^{\prime}$ coupling to run much more quickly than usual to very small values as the scale under consideration is lowered. In other words, the effective anomaly provides the remainder of the qualitative justification for the very small values of $g_{Z^{\prime}}$ and $\sin\phi_{Z}$ of this model.
At a scale $\mu$ below the unification scale but above $\tilde{\nu}_{M}$, the running coupling constants for the SU(N) groups obey the following equation:
$\frac{4\pi}{g_{N}^{2}(\mu)}=\frac{4\pi}{g_{N}^{2}(M_{U})}-\frac{b_{N}}{2\pi}\ln\left(\frac{\mu}{M_{U}}\right)$ (4.2)
$b_{N}=\left(-\frac{11}{3}N+\frac{1}{3}n_{f}+\frac{1}{6}n_{s}+\frac{2}{3}Nn_{fA}+\frac{1}{3}Nn_{sA}\right),$
where $n_{f}$ and $n_{s}$, $n_{fA}$ and $n_{sA}$ are the numbers of fundamental fermion and scalar N-tuplets, and the numbers of fermion and scalar adjoint representations, respectively. In all cases, the fermions are 2-component Weyl fermions and the scalars are complex.
In this model, $n_{fA}=0$ and $n_{sA}=1$ for each SU(N) group. The particle content to use between $M_{U}$ and $\tilde{M}_{\nu}$ is the following:
  1. 1.
    Up quarks: 4 W triplets and 2 non-W triplets
  2. 2.
    Down quarks: 4 W triplets and 4 non-W triplets
  3. 3.
    Charged leptons: 6 W and 2 non-W
  4. 4.
    Neutrinos: 6 W and 4 non-W
  5. 5.
    SU(3) scalars: 6 triplets
  6. 6.
    SU(2) scalars: 6 doublets (2x flavors 1 & 2, 2x adjoint)
  7. 7.
    Singlet scalars: 6 (2x flavors 1 & 2, 2x adjoint),
where in the notation of this paper, a “W” fermion is part of an isodoublet that interacts with the W boson.
The scalars fall into these categories for the following reasons: The $m=2$ flavor 3 triplets get eaten (or made massive) by the unification-scale symmetry breaking. Of the 9 components of the $m=2$ U(3) adjoint scalars, 3 are an SU(2) adjoint multiplet, 4 form 2 doublets, and the remaining 2 form 2 singlets.
With the above particle content, the beta factors for running of couplings between $M_{U}$ and $\tilde{M}_{\nu}$ are:
$b_{3}$ $=\left(-\frac{11}{3}3+\frac{1}{3}14+\frac{1}{6}6+\frac{1}{3}3\right)=-\frac{13}{3}$ (4.4)
$b_{2}$ $=\left(-\frac{11}{3}2+\frac{1}{3}18+\frac{1}{6}6+\frac{1}{3}2\right)=\frac{1}{3}.$
Between the $\tilde{M}_{\nu}$ scale and the electroweak scale, $b_{3}$ remains the same, but $b_{2}$ is reduced since the number of massless lepton doublets (and corresponding scalar doublets) is reduced by 1. So in this mass range $b_{2}=-1/6$.
In the rough approximation that the $\tilde{M}_{\nu}$ scale is one third of the way between the electroweak and unification scales, an approximate $b_{2}$ to use all the way from the electroweak scale to the unification scale is
$b_{2}\simeq\frac{1}{6}\textrm{ for }M_{Z}\textrm{ to }M_{U}.$ (4.5)
Using the measured values of the SU(2) and SU(3) coupling constants at the scale of $M_{Z}$, the unification scale is determined to be
$M_{U}\simeq 6\times 10^{14}\textrm{ GeV}.$ (4.6)
The inverse of the nonAbelian coupling at the unification scale is:
$\alpha_{2}^{-1}\left(M_{U}\right)=\alpha_{3}^{-1}\left(M_{U}\right)\simeq 29,$ (4.7)
[…]
In this model, the beta factor for the weak hypercharge coupling $g_{Y}$ is approximately given by
$b_{Y}\simeq$ (4.8)
$\simeq\frac{3}{20}\left(2\left(\left({\tfrac{1}{3}}\right)^{2}8+\left({\tfrac{2}{3}}\right)^{2}4+\left({\tfrac{4}{3}}\right)^{2}2\right)+\left(\left({\tfrac{2}{3}}\right)^{2}4+\left({\tfrac{4}{3}}\right)^{2}2\right)\right)$
$+\frac{3}{20}\left({\tfrac{2}{3}}\left(\left(1\right)^{2}6+\left(1\right)^{2}6+\left(2\right)^{2}2\right)+{\tfrac{1}{3}}\left(\left(1\right)^{2}12+\left(2\right)^{2}2\right)\right)$
$b_{Y}\simeq\frac{17}{3}$
where the non-W neutrinos and the flavor 2,3 singlet scalars all have zero hypercharge.
Assuming that the $Y$ coupling runs normally and is not affected by the heavy-neutrino-induced effective anomaly, the hypercharge coupling at the unification scale would be
$\alpha_{Y}^{-1}\left(M_{U}\right)\simeq 32.$ (4.9)
However, just as for the $Y^{\prime}$, it is possible that $g_{Y}$ runs down a little faster from $\tilde{M}_{\nu}$ to $M_{Z}$ (or runs up a little faster from $M_{Z}$ to $\tilde{M}_{\nu}$). It is proposed that if this effect was fully taken into account, $g_{Y}$ may unify with the nonAbelian couplings at the unification scale, having $\alpha_{Y}^{-1}\left(M_{U}\right)\simeq 29$.
As mentioned above, $g_{Y^{\prime}}$ also runs more quickly below the scale of $\tilde{M}_{\nu}$ due to a combination of a larger beta function and the heavy neutrino-induced effective anomaly. It is proposed that if the effect of the effective anomaly was fully taken into account, a $g_{Y^{\prime}}$ that unified with the other couplings at the unification scale could potentially run down fast enough (transforming to $g_{Z^{\prime}}$ at $M_{Z}$) in order to reach the extremely small values required to reproduce the measurements of $Z^{\prime}$ boson interactions described in section 5. Future work would be required to determine whether the suggested scenario can be supported by detailed calculations.
5 Experimental Implications
The model proposed in this paper is very different than the Standard Model. To truly define this model, more detailed calculations would need to be performed that are outside the scope of this paper. But even in the absence of these calculations, a number of statements can be made about features (e.g. masses, couplings) the model would have to have in order to reproduce experimental data.
[…]
5.1 A Light Z’ Boson
A consequence of the structure of this model is the existence of an additional light U(1) gauge boson – a $Z^{\prime}$. In order to be consistent with precision electroweak experiments, the angle $\phi_{Z}$ (from eqs (2.16) and (2.17)) that determines the mass and coupling of the $Z^{\prime}$ must be very small. In section 4, a mechanism was suggested for this very small value. The calculations in this paper are not detailed enough to predict the mass and coupling of the $Z^{\prime}$ boson. Instead, a phenomenological approach is taken; an observed experimental anomaly is studied and mapped to the $Z^{\prime}$ of this model.
[…]
At the electroweak scale, the generalized Weinberg rotation of eq (2.16) rotates the $T_{2}^{3}$, $T^{Y}$, and $T^{Y^{\prime}}$ matrices of eqs (1.7) and (2.15) to $T^{\gamma}$, $T^{Z}$, and $T^{Z^{\prime}}$. Using the angles from eq (2.17), the $Z^{\prime}$ matrices take the form:
$g_{Z^{\prime}}T_{1}^{Z^{\prime}}$ $=-{\rm{diag}}\left(1+\tfrac{2}{3}x,1+\tfrac{2}{3}x,1+\tfrac{2}{3}x,0,x-1,x-1\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$ (5.1)
$g_{Z^{\prime}}T_{2}^{Z^{\prime}}$ $=-{\rm{diag}}\left(\tfrac{1}{3}-\tfrac{1}{3}x,\tfrac{1}{3}-\tfrac{1}{3}x,\tfrac{1}{3}-\tfrac{1}{3}x,\tfrac{1}{6}-x,-\tfrac{5}{6},-\tfrac{5}{6}\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$
$g_{Z^{\prime}}T_{3}^{Z^{\prime}}$ $={\rm{diag}}\left(\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,x-1,0,0\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$
$g_{Z^{\prime}}T_{A}^{Z^{\prime}}$ $={\rm{diag}}\left(\tfrac{31}{12}+\tfrac{1}{3}x,\tfrac{31}{12}+\tfrac{1}{3}x,\tfrac{31}{12}+\tfrac{1}{3}x,-\tfrac{37}{12}+x,-\tfrac{25}{12},-\tfrac{25}{12}\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}\,,$
where the last line is the group structure of $Z^{\prime}$ acting on adjoint fields, and $x=\sin^{2}\theta_{Z}\simeq\sin^{2}\theta_{W}$ for small $\sin\phi_{Z}$. The group structure of the $Z^{\prime}$ acting on gauginos is the same as that of $T_{2}^{Z^{\prime}}$.
Given those group structures and the definitions of quarks and leptons earlier in the paper, the following couplings with the $Z^{\prime}$ boson can be read off of eq (5.1):
$\begin{array}[]{|r|c|c|}\hline\cr&g^{L}&g^{R}\\ \hline\cr u&-1-\tfrac{2}{3}x&-\tfrac{1}{6}-\tfrac{2}{3}x\\ d&\tfrac{14}{3}+\tfrac{1}{3}x&-\tfrac{7}{6}+\tfrac{1}{3}x\\ e&1-x&1-x\\ \nu_{e}&0&0\\ \hline\cr\end{array}\,\,\,\times\frac{e\sin\phi_{Z}}{\sin\theta_{W}\cos\theta_{W}}$ (5.2)
If $\sin\phi_{Z}$ takes the following value
$\frac{e\sin\phi_{Z}}{\sin\theta_{W}\cos\theta_{W}}\simeq 3.4\times 10^{-4}\,,$ (5.3)
then the quark couplings can be translated into Vector (V) and Axial-vector (A) proton, neutron and electron couplings as follows:
$\begin{array}[]{|r|c|c|}\hline\cr&V&A\\ \hline\cr p&C_{p}\simeq 0.12&a_{p}\simeq 0.49\\ n&C_{n}\simeq 1.0&a_{n}\simeq-0.95\\ e&|\epsilon|\simeq 0.87&0\\ \hline\cr\end{array}\,\,\,\times 10^{-3}\,,$ (5.4)
where notation from [37] (Appendix C) is being used for protons and neutrons, and $\epsilon$ is the electron vector coupling divided by $e\simeq 0.3$.
[…]
The mass of the $X17$ in this model is generated by its interaction with the scalar field $\tilde{\phi}_{22}$, and its vev $\bar{\tilde{\phi}}_{22}$. From the form of $g_{Z^{\prime}}T_{2}^{Z^{\prime}}$ in eq (5.1), the tree-level mass is given by:
$m^{2}_{Z^{\prime}}\simeq\bar{\tilde{\phi}}_{22}^{2}\left(\tfrac{5}{6}\right)^{2}\left(\frac{e\sin\phi_{Z}}{\sin\theta_{Z}\cos\theta_{Z}}\right)^{2}\,.$ (5.5)
From the quark mass matrices, the vev $\bar{\tilde{\phi}}_{22}$ should be only a couple of GeV. If so, then the tree-level $Z^{\prime}$ mass is only about 2 MeV. It is assumed that the remainder of its 17 MeV mass comes from radiative corrections.
5.2 A Seventh Quark
This model predicts the existence of a seventh flavor of quark. The model itself does not predict the mass of the quark, but if the mass was larger than that of the top quark, then the seventh quark could potentially break electroweak symmetry at its mass scale, not that of the top quark. In other words, the model appears to prefer a seventh quark with a mass smaller than that of the top quark.
A recent paper shows that if there is a fourth down-type quark with a mass of 2.8 GeV, then most of the exotic hadrons discovered over the last twenty years fit nicely into the quark model as normal mesons and baryons involving this quark, rather than 4- or 5-quark hadrons [40].
If a fourth down-type quark existed, one might expect it to generate predictions for CKM data that no longer agree with experimental data. The opposite is true. The CKM matrices of eq (3.22) accurately reproduce absolute values of CKM element data, even for some data that disagree with the Standard Model by $\sim 3\sigma$. In particular, the model reproduces data that cast doubt on CKM unitarity and on the W boson only connecting left-handed quarks [41].
A natural question is how an additional 2.8 GeV quark could have evaded detection so far. Most of the dozens of direct searches for a new down-type quark surprisingly do not rule out a quark of the type proposed here [42]. In particular, most model-independent searches do not exclude a new quark with a mass smaller than that of the beauty quark.
One notable exception is the fact that inclusive hadronic cross section data seem to rule out the possibility of a light seventh quark. The Standard Model reproduces these data very well, apparently leaving no room for an additional quark. However, it was recently proposed that an assumption inherent in Standard Model calculations may be incorrect [43]. Namely, calculations assume that there is no interference between $Z/\gamma^{*}\to u\bar{u}\to X$ and $Z/\gamma^{*}\to d\bar{d}\to X$ events, where $X$ is the same hadronic state in both processes (e.g. $\pi^{+}\pi^{-}$). If interference is taken into account, the Standard Model may underpredict inclusive hadronic cross section data below the $Z$ pole. If so, the additional quark may be required in order to fit the experimental data.
Another difficulty with the model is the fact that the charm quark of this model has a different coupling to the $Z$ boson than in the Standard Model. In [41], a mechanism is proposed whereby the effects of the different charm coupling combined with the additional quark decaying to charm allow this model to reproduce the LEP-measured charm cross section and asymmetry data.
But the mechanism described there would not help reconciling $Z$ boson width measurements with the larger total coupling of the charm quark in this model and the presence of an additional quark. A mechanism that could allow this model to reproduce those data is described in the $Z$ width section below.
Under the hypothesis that the seventh quark exists and has a mass of around 2.8 GeV, a large number of predictions have been made for new exotic hadrons that could be observed in current experiments [40].
5.3 Right-Handed Neutrinos
In the neutrino part of section 3, it was proposed that $\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$ are right-chiral neutrinos that are isodoublets with the right-chiral muon and tau lepton, respectively. These right-chiral neutrinos have Dirac mass terms together with two of the model’s isosinglet neutrinos ($\tilde{\nu}^{(2)}$ and $\tilde{\nu}^{(3)}$). That makes the heavy right-chiral neutrinos of this model different than most models that form the basis for experimental searches. In those models, the heavy right-chiral neutrino is usually an isosinglet that has a Dirac mass term mixing it with one of the observed left-chiral isodoublet neutrinos. This model difference means that some of the heavy neutrino searches do not set limits on this model.
With that background, it is assumed that the masses of $\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$ are an MeV or so heavier than the mass of the tau lepton. Despite many searches that produced no evidence for heavy neutral leptons, it appears that neutrinos of the masses and type proposed here may not be ruled out.
Figure 4 of [44] shows exclusion limits for a massive neutrino that interacts with the tau lepton. The only existing (as opposed to projected) experiment shown in that figure pertinent to the mass of the heavy tau neutrino proposed above is DELPHI. The DELPHI experiment searched for $Z\to\nu_{h}\bar{\nu}$ events, where $\nu_{h}$ is a heavy neutrino and $\bar{\nu}$ is a normal (almost massless) antineutrino [45]. In the model proposed here, the $Z$ boson has no direct coupling of the heavy right-chiral $\nu_{W}^{(3)}$ tau neutrino with the normal neutrinos $\nu_{W}^{(1)}$, $\tilde{\nu}_{W}^{(2)}$ and $\tilde{\nu}_{W}^{(3)}$. Also, the model’s mass parameters can accommodate heavy-light neutrino mass mixing that is below the DELPHI limit. In the simplified example matrix of eq (3.27), that would require choosing parameters such that $m_{133}<.01m_{\nu 3}$. In this way, the model can evade the DELPHI limits for a heavy tau neutrino.
Figure 1 of [46] shows exclusion limits for a massive neutrino that interacts with the muon. Experiments performed by DELPHI, L3 [47], CHARM II [48] and NA3 [49] limit the mass. L3 looked for heavy neutrinos in the same way as DELPHI (searching for $Z\to\nu_{h}\bar{\nu}$ events). CHARM II searched for transitions from normal muon neutrinos to heavy ones in neutral current scattering. For $m_{122}<.005m_{\nu 2}$, the right-chiral muon neutrino of this model evades the DELPHI, L3 and CHARM II limits. NA3 was not sensitive to a heavy neutrino as short-lived as the one considered here. So the heavy right-chiral muon neutrino considered here is not excluded by the experiments quoted in [46].
A CMS search for heavy neutrinos from $W$ decay only considered squared interaction strengths up to 0.5 [50]. Since the neutrino considered here has an interaction strength of 1, it is not excluded.
Despite its short life, the CMS search from $B$ meson decay [51] could potentially have seen some signal of the heavy muon neutrino considered here. That experiment ruled out a neutrino of this type, unless its mass was in their veto range of 1.74 to 1.8 GeV. The reason for the veto is because that range could be contaminated by an unknown number of $D^{0}\to K\pi$ decays. The CMS data did show a feature in that range, but it was attributed to $D^{0}$ decays and excluded from the analysis. It is assumed that some part of that feature was due to the heavy muon neutrino proposed here.
Neutrinos of the type discussed here are too heavy to be produced in the decay of the $D$ meson or tau lepton, so their presence would not contradict $D$ or tau decay branching ratios or the fact that measured Michel parameters are consistent with only left-chiral leptons being produced in decays of tau leptons [52, 53].
The heavy neutrinos proposed here are not sterile; they interact with the $Z$ boson. Therefore, they add to the inclusive cross sections that were measured very precisely near the $Z$ peak at LEP. A way to reconcile this model with those measurements is discussed in the $Z$ width section below.
As noted in [5], experimentally observed branching ratios for $B\to D^{(*)}\tau\nu$ exceed Standard Model predictions by $>3\sigma$. Also, the number of observed multilepton LHC events exceeds Standard Model predictions by $>5\sigma$. This model provides the following new physics explanation for those anomalies.
Heavy neutrinos of the type proposed here could be produced in the decay $B\to D^{(*)}\mu\nu_{W}^{(2)}$ with $\nu_{W}^{(2)}$ then decaying to two charged particles (accompanied by a normal neutrino for pure leptonic decays). Since the mass of $\nu_{W}^{(2)}$ is similar to that of the tau lepton, these decays could look like the decay $B\to D^{(*)}\tau\nu$, with $\tau$ decaying to three charged particles. If an experiment grouped these decays that way (or grouped some of them that way), then the experimental results would show an apparent branching ratio for $B\to D^{(*)}\tau\nu$ decays that was significantly larger than the branching ratio predicted by the Standard Model. It is proposed that this is the origin of the $B\to D^{(*)}\tau\nu$ anomaly.
In LHC events, the heavy neutrinos of this model lead to many more multilepton events. As an example, it is interesting to consider a $t\bar{t}$ event, where the $t$ decays to a b-jet and a $W^{+}$ boson. In the Standard Model, there is some probability that the $W^{+}$ produces a charged lepton and a neutrino. In this model, that probability is doubled for production of muons and tau leptons, due to production involving their heavy neutrinos. Those heavy neutrinos also decay to at least one additional charged lepton. The net result is that this model predicts many more multilepton events at LHC than does the Standard Model.
5.4 Muon and Tau Asymmetry Data
The muon and tau lepton of this model have different couplings to the Z boson than in the Standard Model. Those differences must be reconciled with the fact that cross section and asymmetry data from $e^{+}e^{-}\to Z/\gamma^{*}\to\mu^{+}\mu^{-}$ or $\tau^{+}\tau^{-}$ decays are consistent with the muon and tau lepton having the same coupling to the Z boson as in the Standard Model [54].
The couplings to the Z boson for the left- and right-chiral muons and tau leptons for the Standard Model (SM) and this model (Model) are:
$\begin{array}[]{|r|c|c|}\hline\cr\textrm{Z coupling}&g^{L}_{\mu,\tau}&g^{R}_{\mu,\tau}\\ \hline\cr\mu,\tau\textrm{ (SM)}&-\tfrac{1}{2}+\sin^{2}\theta_{W}&\sin^{2}\theta_{W}\\ \mu,\tau\textrm{ (Model)}&-\tfrac{1}{2}+\sin^{2}\theta_{W}&-\tfrac{1}{2}+\sin^{2}\theta_{W}\\ \hline\cr\end{array}\,.$ (5.6)
It can be seen that if some of this model’s new physics was able to modify the effective right-chiral Z-couplings of the muon and tau lepton by $\delta g^{R}_{\mu,\tau}=\tfrac{1}{2}$, then their effective Z-couplings would be the same as those in the Standard Model, and consistent with the data.
From eqs (2.18) and (2.20), it can be seen that the right-handed muon and tau neutrinos discussed in the last section ($\nu_{W}^{(2)}$ and $\nu_{W}^{(3)}$) have couplings with the Z boson of $\tfrac{1}{2}$.
In this model, there would have been events at LEP where a $Z$ boson decayed to $\nu_{W}^{(3)}\nu_{W}^{(3){\dagger}}$. If the mass of $\nu_{W}^{(3)}$ is only slightly larger than the tau lepton, then it will decay by $\nu_{W}^{(3)}\to\tau^{-}e^{+}\nu_{e}$. Therefore, $Z\to\nu_{W}^{(3)}\nu_{W}^{(3){\dagger}}$ events would have looked like $Z\to\tau^{-}\tau^{+}e^{+}e^{-}$ events. If the experiments considered $e^{+}e^{-}$ to be radiated pairs, then these events would have been grouped together with the remaining $Z\to\tau^{-}\tau^{+}$ events. In that case, the effective right-chiral tau lepton coupling produced by this model would have been very similar to that of the Standard Model and therefore would have reproduced the data.
Similarly, decays of $Z\to\nu_{W}^{(2)}\nu_{W}^{(2){\dagger}}$ may have been grouped together with $Z\to\mu^{-}\mu^{+}$ events. Again, the effective right-chiral muon coupling of this model would then have been very similar to that of the Standard Model.
There is a subtlety with this scenario. If 4-lepton events had been rejected by the tagging at LEP or SLD, then a significant number of right-chiral neutrino events would have been excluded from effective coupling data. However, it appears that these events were not excluded. It was stated that “The classification of final states with radiated fermion pairs, i.e. of four-fermion events, into one of the three lepton categories is made by choosing the lepton pair with the highest invariant mass” [54].
It is also interesting that about 1.2% of $e^{+}e^{-}\to\mu^{+}\mu^{-}$ events observed by OPAL at LEP had two muons with the same charge. It was mentioned that “the origin of this effect is not understood” [55], and it was assumed that one of the muons had been misidentified. An alternative explanation in this model is provided by events involving the decay $\nu_{W}^{(2)}\to\mu^{-}\mu^{+}\nu_{\mu}$. Rather than pair the $\mu^{-}$ from $\nu_{W}^{(2)}$ with the $\mu^{+}$ from $\nu_{W}^{(2){\dagger}}$, some events could have paired the $\mu^{+}$ from $\nu_{W}^{(2)}$ with the $\mu^{+}$ from $\nu_{W}^{(2){\dagger}}$.
In summary, it is proposed that the decay of Z bosons to right-chiral neutrinos that were then tagged as right-chiral muons and tau leptons could have allowed the model presented here to have reproduced the muon and tau lepton cross section and asymmetry measurements at LEP and SLD.
5.5 The Z boson width
Precision experiments performed at LEP and SLD established the width of the $Z$ boson and how much of that width was due to the $Z$ boson decaying to hadrons, charged leptons, and invisible particles (light neutrinos) [54]. The model presented here has additional Beyond-Standard-Model particles that would have also been produced in $Z$ boson decay: an additional down-type quark and two additional neutrinos with masses similar to that of the tau lepton. The model also has $Z$ interactions with the right-chiral muon, tau lepton and charm quark that are larger than in the Standard Model. This subsection qualitatively presents a way that this model could potentially reproduce the $Z$ boson width data.
The LEP hadronic cross sections were calculated by counting hadronic events (less background) and dividing by the luminosity (and efficiency). Leptonic cross sections were calculated in a similar way. If it is assumed that the actual luminosity of these events was $\sim$13lower than the luminosity used in the analysis of the LEP experiments, then it would be possible for this model to reproduce the LEP $Z$ boson width data (along with the various ratios of hadronic cross sections to leptonic cross sections).
At LEP, the luminosity was estimated by measuring the number of low angle $e^{+}e^{-}\to e^{+}e^{-}$ events (angles of $\sim 25-100$ milliradians relative to the beam line). If the Standard Model is assumed, then these events are almost entirely due to t-channel photon exchange Bhabha scattering. Since only QED is involved, this process can be calculated with high accuracy, leading to very precise luminosity estimates (see for example [56]).
In this model, there are other processes that contribute to low angle $e^{+}e^{-}\to e^{+}e^{-}$ events. First, a t-channel diagram involving the $Z^{\prime}$ boson of this model can contribute, and its diagram interferes constructively with the photon exchange diagram. Approximating the $Z^{\prime}$ boson as massless, from eq (5.1), the t-channel Bhabha scattering cross section in this model compared to that of the Standard Model is:
$\frac{\sigma_{Bh}}{\sigma^{SM}_{Bh}}\simeq\left(1+\frac{\sin^{2}\phi_{Z}}{\tan^{2}\theta_{W}}\right)^{2}\,.$ (5.7)
A value of $\sin\phi_{Z}\simeq 0.1$ at the $Z$ boson mass scale would translate into about a 7% increase in t-channel $e^{+}e^{-}\to e^{+}e^{-}$ events, and a luminosity about 6% smaller than calculated using the Standard Model.
Another process could potentially have an even larger impact. In section 3, it was proposed that the electron’s mass is generated by its interaction with the adjoint scalar $(\varphi_{2})_{32}$ that has a vev of $\bar{\varphi}_{2}$. As discussed in the scalar section below, the mass of this scalar could be small, potentially even smaller than that of the electron.
If $(\varphi_{2})_{32}$ does have a small mass, the t-channel exchange of this scalar could also generate more low angle $e^{+}e^{-}\to e^{+}e^{-}$ events for the same beam luminosity than the Standard-Model prediction. Between the $Z^{\prime}$ and this scalar, the model could produce $\sim$15more t-channel $e^{+}e^{-}\to e^{+}e^{-}$ events than the Standard Model. That would imply that a luminosity calculated using this model would be $\sim$13lower than the luminosity calculated using the Standard Model.
If that were the case, then the number of hadronic events observed would correspond to a $\sim$15larger hadronic cross section than predicted by Standard-Model calculations, so Standard-Model-only $Z$ boson decays would not be able to reproduce the data. The same is true of leptonic decays. The model presented here, however, could reproduce the observed cross sections due to the additional particles and different $Z$ couplings mentioned at the beginning of this section.
The $(\varphi_{2})_{32}$ scalar also interacts with the $Z$ boson, so it would be produced in pairs in $Z$ boson decays. Due to its low mass and zero charge, $(\varphi_{2})_{32}$ would add to the invisible width of the $Z$ boson.
But this scalar would also contribute to the initial state radiation that significantly increases the actual width of the $Z$ boson compared to what is observed (see for example fig. 1.12 of [54]). That additional increase in width could potentially accommodate the additional invisible partial width of $(\varphi_{2})_{32}$.
To end this section, it is worth pointing out that the example value of $\sin\phi_{Z}$ mentioned above corresponds to $\cos\phi_{Z}\simeq 0.995$. From eq (3.6), and given the extraordinary precision in measurements of $\sin^{2}\theta_{W}$, such a value would require significant new physics quantum contributions to the $W$ boson mass. This model certainly has a lot of new physics that could make those contributions, and their effect would need to be checked to see if a consistent picture can be achieved.
An alternative scenario is that $\sin\phi_{Z}$ is much smaller at the electroweak scale, so no (or very small) new physics contributions would be needed to reproduce precision $W$ mass measurements. In that scenario, the entire luminosity correction described above is attributed to the $(\varphi_{2})_{32}$ scalar.
5.6 The Muon Magnetic Moment
The main source of uncertainty in a Standard-Model calculation of the $g-2$ magnetic moment of the muon comes from calculation of loop diagrams that involve hadrons. There are two main competing ways to calculate the hadronic contribution: (i) a data-driven approach [57, 58, 59] and (ii) a lattice QCD approach [60].
This model has both an additional quark and also quark couplings that are different from those in the Standard Model. Therefore, a lattice calculation using this model would produce a different result than one using the Standard Model. On the other hand, a data-driven approach based on this model should produce similar results to one based on the Standard Model. The normal data-driven approach assumes that all Standard-Model nonperturbative QCD interactions are incorporated into the low energy data used as input. Changing the model just means changing that assumption (replacing Standard-Model with this model). But since the data are still the same, the hadronic contribution to the calculation would remain the same.
As a result, since no lattice calculations involving this model are available, this model’s $g-2$ results should be compared to those of the data-driven calculation. Also, any differences from the Standard Model result should have their origin in non-hadronic parts of the calculation.
Experimental measurement of the anomalous magnetic moment of the muon produces results that differ from data-driven Standard Model (SM) calculations by 5.1 standard deviations [57, 61, 59]:
$a_{\mu}^{\rm{exp}}-a_{\mu}^{\rm{SM}}=249\pm 48\times 10^{-11}.$ (5.8)
This model may be able to reproduce that value as a result of the model’s differences in the weak sector (as opposed to the hadronic sector).
In the Standard Model, the one-loop contribution to the muon magnetic moment involving the W and Z bosons is [58, 62, 63]:
$a_{\mu}^{\rm{EW}}[\rm{1loop}]=\frac{G_{F}m_{\mu}^{2}}{8\sqrt{2}\pi^{2}}\left(\tfrac{7}{3}+1+\tfrac{4}{3}g_{\mu V}^{2}-\tfrac{20}{3}g_{\mu A}^{2}\right).$ (5.9)
The four terms are from diagrams involving (i) two $W$ bosons, (ii) one $W$ boson and one Goldstone-boson component of the Higgs doublet (fig. 100c of [58]), (iii) a vector current $Z$ boson connection, and (iv) an axial vector $Z$ boson connection. Plugging in the Standard Model couplings, one gets $a_{\mu}^{\rm{EW}}[\rm{1loop}]\simeq 194.8\times 10^{-11}$ [62].
This model’s muon only has a vector-current connection to the $Z$ boson, and it is larger than in the Standard Model. Consequently, the last term of eq (5.9) vanishes, and the third term is larger than in the Standard Model. This model’s muon does not have a tree-level coupling with the Higgs doublet (since the muon does not get its mass from the Higgs boson), so the second term vanishes (or is very small). On the other hand, the right-chiral muon neutrino of this model causes the first term above to get doubled. After plugging in the couplings from eq (5.6), this model generates $a_{\mu}^{\rm{EW}}[\rm{1loop}]\simeq 549\times 10^{-11}$
In the Standard Model, the 2-loop calculation reduces 194.8 down to 153.6 [62], and the 3-loop contribution is negligible. In other words, consideration of 3 loops reduces the 1-loop result by 21%.
Making the very rough assumption that a 3-loop calculation in this model would also reduce the 1-loop result by around 21%, this model’s 3-loop value would be in the range of $433\times 10^{-11}$. If that is compared to the Standard Model value of 153.6, the result is
$a_{\mu}^{\rm{model}}-a_{\mu}^{\rm{SM}}\sim 279\times 10^{-11}.$ (5.10)
[…]
5.7 A Fourth Charged Lepton
This model predicts the existence of a fourth charged lepton (referred to in this paper as the “Omega lepton”). As mentioned in section 3, direct searches have ruled out a fourth charged lepton with a mass of less than 103 GeV [64]. If the Omega lepton had a mass of around 2 TeV, it could provide an explanation for hints of new particles seen by the CMS experiment [65].
In the appendix, it is argued that the quantum-generated superpotential causes the complex scalar fields $(\tilde{\phi}_{21})_{2}$ and $(\phi_{21})_{3}$ to have masses similar to (or a bit larger than) that of the Omega lepton (where subscripts outside of parentheses denote the $m=2$ SU(3) index of those scalar triplets). The gauge interaction of the $(\tilde{\phi}_{21})_{2}$ scalar includes a term that connects it to a W boson, a Z boson, and the vev $\bar{\tilde{\phi}}_{21}$. This vev would allow the $(\tilde{\phi}_{21})_{2}$ charged scalar to decay into a W and a Z boson.
[…]
5.8 Scalars
The model includes the following complex scalars: 12 triplets, 2 octets, and 2 singlets. This section describes a proposed scenario for masses of those scalars.
[…]
As discussed in section 3, in an $N=2$ Super Yang-Mills (SYM) theory where the adjoint superfield has a tree-level mass, the adjoint scalars become massless color monopoles. The monopoles condense and cause confinement. Since the $m=1$ sector of this theory has a very similar moduli space, it was proposed in section 3 that the octet of $m=1$ SU(3) adjoint scalars of this theory become monopoles, condense, and cause color confinement in the same way.
The six $m=1$ scalar triplets are leptoquarks since they connect gaugino quarks with leptons (and possibly also adjoint-representation quarks with leptons). The model puts no constraint on the scalar masses $m_{1F}$ and $\tilde{m}_{1F}$. It is therefore assumed that these masses are larger than the current lower limits from leptoquark searches. Having large masses for these scalars also strengthens the above confinement argument, since an effective theory below the leptoquark masses has an $m=1$ moduli space that is even more similar to that of $N=2$ SYM. If the flavor-1 leptoquarks have masses of around 10 TeV, they could be at least partly responsible for the non-resonant di-electron anomaly discussed in [5].
The model requires the masses $\tilde{m}_{11}$ and $m_{22}$ of the scalar triplets $\tilde{\phi}_{21}$ and $\phi_{22}$ to be larger than the electroweak scale. The neutral scalar $(\phi_{22})_{2}$ has the right structure from its gauge interactions to generate di-photon, di-Z, di-W, and $b\bar{b}$ events. If $(\phi_{22})_{2}$ had a mass of $\sim$650 GeV it could be responsible for the excesses seen in these events. Quantum effects could generate larger masses of $\sim$850-950 GeV for the charged and neutral scalars $(\phi_{22})_{1}$ and $(\phi_{22})_{3}$. These scalars could decay into two jets and be responsible for the di-jet excesses seen at these masses by CMS and ATLAS.
It was suggested above in the fourth charged lepton section that the charged $(\tilde{\phi}_{21})_{2}$ scalar may have a mass of 2.1 or 2.9 TeV. It is possible that the neutral and charged scalars $(\tilde{\phi}_{21})_{1}$ and $(\tilde{\phi}_{21})_{3}$ may have masses of $\sim$3.5 TeV. From the quantum-generated superpotential coupling of $(\tilde{\phi}_{21})_{1}(f_{22})_{2}(\phi_{22})_{3}$ and the vev $\bar{\tilde{\phi}}_{21}$, the $(\tilde{\phi}_{21})_{1}$ scalar could decay to $(\phi_{22})_{3}(\phi_{22})_{3}^{\dagger}$. It could then be responsible for the di-di-jet excess seen at the LHC. Quantum couplings could also cause the charged scalar $(\tilde{\phi}_{21})_{3}$ to decay to $tb$, generating the excess seen in those data. The anomalies described in this paragraph and the one before it are discussed (and cited) in [5].
The observed Higgs Boson accounts for the SU(2) doublet within the triplet $\phi_{21}$. It was proposed in the fourth lepton section above that the same mechanism that gives a large mass to the Omega lepton also gives a similar mass to the remaining scalar in the triplet: $(\phi_{21})_{3}$.
The $W$-scale mass from eq (2.8) together with radiative corrections could generate a mass of around 150 GeV for scalars in the $\tilde{\phi}_{22}$ triplet. Similarly, quantum effects could generate masses of around 95 GeV for the SU(2) adjoint triplet scalars within the $\varphi_{2}$ octet. If so (or if the masses were reversed), these scalars could potentially be responsible for the remaining Higgs-like resonant signals discussed in [5].
A superpotential coupling $\tilde{\Gamma}_{211}$ together with the Higgs vev $\bar{\phi}_{21}$, generates a mass for the charged scalar $(\varphi_{2})_{31}$. This mass is assumed to be larger than that of the $W$ boson. Similarly, a large mass is assumed to be generated for the charged scalar: $(\varphi_{2})_{13}$.
The scalars that have not yet been discussed in this section are $(\varphi_{2})_{32}$, $(\varphi_{2})_{23}$ the scalar in $\varphi_{2}$ proportional to $t^{8}$ and the two adjoint singlet scalars. It is assumed that the first of these has a small mass. It was discussed above how this small mass could help this model to reproduce experimental $Z$ boson width measurements. The last three scalars mentioned above do not interact with any gauge bosons below the unification scale. One combination of these scalars ($(\varphi_{2})_{33}$) also has no (or very little) interaction with any sub-unification-scale fermions or scalars. For that reason, it is an interesting dark matter candidate.
Discussion
The theory presented in this paper is being proposed as an alternative to the Standard Model. The paper has taken a two-pronged approach: theoretical and phenomenological.
In sections 1, 2, 4 and the appendix, a model is presented that has a number of attractive theoretical features. For example, it is holomorphic, invariant to local superspace gauge transformations, supports coupling constant unification, and is similar at the unification scale to a theory that has been shown to be free of quadratic divergences to at least two loops.
[…]
The theory looks promising on the theoretical side and also on the phenomenological side. More work needs to be done to tie these two sides together. Would actual nonperturbative calculations support the parameter values required to reproduce data?
But even without that work, the model provides a couple of interesting explanations and makes a number of predictions. The model provides explanations for the mechanisms of confinement and neutrino oscillations. The model predicts a fourth charged lepton, a seventh quark (without an eighth), a $Z^{\prime}$ boson, right-chiral neutrinos with masses of a few GeV, right-chiral quark interactions with the W boson, and dozens of additional scalar particles. So far, these predictions do not appear to be ruled out by existing data; in fact, they could provide new physics explanations for many of the anomalies discussed in [5].
Appendix A Quantum Effective potential
This appendix proposes possible superpotential terms that may be generated by nonperturbative quantum effects.
In [17, 18, 14, 66, 67, 68], it is shown how holomorphy and symmetry arguments can be used to determine the exact superpotential terms that get generated nonperturbatively for Supersymmetric QCD (SQCD) at low energies. That analysis is anchored in the fact that SQCD is an asymptotically free theory with an ultraviolet renormalization group fixed point.
The U(3)xU(3) theory of this paper has Abelian groups, so it does not appear to be asymptotically free. However, in this appendix it is implicitly assumed that this theory is an effective theory of a more general asymptotically free theory, applicable above the unification scale. With that assumption in hand, this appendix makes arguments similar to those used for SQCD and proposes general features of nonperturbative superpotential terms.
It is possible that instanton calculations could be employed to determine the exact nonperturbative superpotential terms for this model, but that is outside the scope of this paper. Instead, allowed functional forms of terms are derived, and then the magnitude of the quantum-generated couplings that would be required for this model to fit experimental data are phenomenologically identified.
The first step in deriving the effective superpotential terms is to specify the beta function for the two SU(3) groups, evaluated for the case where all tree-level masses and superpotential couplings are zero. The beta function for an SU(N) theory is:
$\beta=g\left(\frac{g^{2}}{16\pi^{2}}\right)b_{N}$ (A.1)
$b_{N}=\left(-\frac{11}{3}N+\frac{1}{3}n_{f}+\frac{1}{6}n_{s}+\frac{2}{3}Nn_{fA}+\frac{1}{3}Nn_{sA}\right),$
where $n_{f}$ and $n_{s}$, $n_{fA}$ and $n_{sA}$ are the numbers of fundamental fermion and scalar N-tuplets, and fermion and scalar adjoint-representation multiplets. In all cases, the fermions are 2-component Weyl fermions and the scalars are complex.
In the model presented in this paper, $n_{f}=18$, $n_{s}=6$, $n_{fA}=0$ and $n_{sA}=1$ for each SU(3) group. The fundamental scalar number comes from 3 flavors of both fundamental and anti-fundamental representations. The fundamental fermions have those plus another 6 fundamental fermion triplets from the gauginos and another 6 from the adjoint superfield. Putting that together, one finds:
$b_{3}=-3.$ (A.2)
Following standard techniques of integrating the one-loop beta function, one finds:
[…]
where $\mu$ is the scale at which the SU(3) coupling is evaluated and $\Lambda$ is the quantum-generated scale of each SU(3) gauge theory.
The next step is to determine which fermion representations generate an axial anomaly. In this model, and using notation similar to that of [13], the axial anomaly is proportional to:
$\sum_{f}\rm{Tr}_{R_{f}}\left(T^{0}_{A_{R}+}\left(\{T^{B}_{V_{R}+},T^{C}_{V_{R}+}\}+\{T^{B}_{V_{R}-},T^{C}_{V_{R}-}\}\right)\right)$ (A.4)
$+$ $\sum_{f}\rm{Tr}_{R_{f}}\left(2T^{0}_{A_{R}-}\{T^{B}_{V_{R}+},T^{C}_{V_{R}-}\}\right).$
In the above expression, $R$ represents the representation of fermion $f$. A representation’s contribution to a local gauge current proportional to $T^{B}_{\pm}$ (as defined in eq (1.7)) is denoted by $T^{B}_{V_{R}\pm}$. Similarly, $T^{0}_{A_{R}\pm}$ denote a representation’s contributions to the global axial currents proportional to $T^{0}_{\pm}$.
Since they are in the (3,3*) representations, the gauginos and the fermions from the twisted adjoint superfield contribute zero to both $T^{0}_{V_{R}+}$ and $T^{0}_{A_{R}+}$. In addition, for the second term of (A.4) they generate a factor of $f^{bad}d^{cad}$ which vanishes. Therefore, those fermions do not contribute to the axial anomaly.
It is speculated that the Abelian gauge field $T^{0}_{V_{R}\pm}$ contributions to the axial anomaly are nontopological and can be ignored, although in [13], it is pointed out that this is not always the case. If they indeed can be ignored, then the remaining contributions to the anomaly for each SU(3) group can be treated independently, and they are identical to those for SQCD with 3 flavors and colors.
In the SQCD derivation of nonperturbative low-energy superpotential terms, the next step is to consider instanton effects. Following that approach, the instanton amplitude in Euclidean space is proportional to:
[…]
where the second equality uses eqs (A.2) and (A.3). For three flavors of fundamental (and anti-fundamental) fields, the instanton amplitude carries an axial U(1) charge of 6. For eq (A.5) to be consistent, the axial U(1) selection rule implies that $\Lambda$ should carry an axial U(1) charge of 2.
The anomaly-generating axial U(1) charges of the superfields $Q_{mF}$, $\tilde{Q}_{mF}$, and $\Phi$ are 1, 1, and 0, respectively. Therefore, the following factors are invariant to the U(3)xU(3) local gauge symmetry and obey the axial selection rule:
${\rm Tr}\left(\Phi^{n}\right)$ (A.6)
$M_{mm^{\prime}FF^{\prime}}^{\left(n\right)}=\Lambda^{-1}\tilde{Q}_{mF}\Phi^{n}Q_{m^{\prime}F^{\prime}},$
where $n$ is a non-negative integer. In the second “meson” configuration above, the flavors are restricted by the local U(1) “+” gauge symmetry to either $F=F^{\prime}=1$ or else $F,F^{\prime}\in\left\{2,3\right\}$.
It should be noted that $M_{mm^{\prime}FF^{\prime}}^{\left(n\right)}$ with $m\neq m^{\prime}$ is proportional to $\theta$, so terms involving these factors do not contribute to the vev of the scalar potential or its minimization, although they do contribute to lepton masses and neutrino mixing as described below.
The following determinants are invariant to the 2/3 global flavor symmetry:
$\Delta_{m_{1}m_{2}m_{3}m_{4}}^{\left(pq\right)}=M_{m_{1}m_{2}22}^{\left(p\right)}M_{m_{3}m_{4}33}^{\left(q\right)}-M_{m_{1}m_{2}23}^{\left(p\right)}M_{m_{3}m_{4}32}^{\left(q\right)}$ (A.7)
Following symmetry arguments similar to those used for SQCD, it is speculated that quantum interactions will generate low energy effective superpotential terms with mass dimension 3 and positive powers of $\Lambda$ that are constructed from the above determinants as well as factors of $M^{\left(n\right)}_{mm^{\prime}11}$, ${\rm Tr}\left(\Phi^{r}\right)$, and other gauge invariant factors such as those mentioned later in this appendix.
An example of a term meeting the above criteria is the following:
$O_{0}=\left({\rm Tr}\left(\Phi^{2}\right)\right)^{2}\left(\sum_{m_{i}}M^{\left(0\right)}_{m_{1}m_{2}11}\Delta_{m_{3}m_{4}m_{5}m_{6}}^{\left(00\right)}\right)^{-1/3}$ (A.8)
where the sum over $m_{i}$ is over all permutations of the $m_{i}$ values being $1$ or $2$. The scalar potential can be derived from the superpotential terms like the one above by taking derivatives to extract “f terms” proportional to $\theta^{2}f$ and inserting them into the f-term part of eq (2.1). The quantum vacuum can be found by minimizing the vev of the scalar potential after including these quantum contributions.
In this process, vevs of scalar potential terms derived from $O_{0}$ will involve terms proportional to
$\bar{O}_{0}=\left\langle O_{0}\right\rangle=\frac{\Lambda\left({\rm tr}\left\langle\varphi_{1}^{2}\right\rangle+{\rm tr}\left\langle\varphi_{2}^{2}\right\rangle\right)^{2}}{\left(\bar{\phi}_{21}\bar{\tilde{\phi}}_{21}\bar{\phi}_{22}\bar{\tilde{\phi}}_{22}\bar{\phi}_{23}\bar{\tilde{\phi}}_{23}\right)^{1/3}}.$ (A.9)
In section 2, it was noted that if the model includes large tree-level masses $\tilde{m}_{21}$ and $m_{22}$, then the vevs $\bar{\tilde{\phi}}_{21}$ and $\bar{\phi}_{22}$ vanish clasically. However, in order to stabilize expressions such as the one above, it is assumed that $\bar{\tilde{\phi}}_{21}$ and $\bar{\phi}_{22}$ acquire small vevs quantum mechanically. This is mentioned in section 3. As shown below, those quantum-generated small vevs lead to the very large right-chiral neutrino Majorana mass that allows the model to reproduce observed neutrino masses and mixing.
The classical scalar potential only restricts the difference $\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$. However, it can be seen from eq (A.9) that the quantum scalar potential will try to make each of these vevs go to infinity while maintaining the difference. A mechanism like that is what causes Supersymmetric QCD with fewer flavors than colors to not have a vacuum solution. But in this theory, a nonzero superpotential coupling $\Gamma_{232}$ can change the result. At tree level, this coupling generates a contribution to the scalar potential of $\Gamma_{232}\bar{\tilde{\phi}}_{23}\bar{\varphi}_{2}\bar{\phi}_{22}$. This contribution vanishes classically, since $\bar{\phi}_{22}$ vanishes classically. But upon quantum generation of $\bar{\phi}_{22}$, this contribution no longer vanishes. Instead, it provides a small counterbalancing effect that stops $\bar{\tilde{\phi}}_{23}^{2}$ from becoming infinite.
It is assumed that minimizing the quantum scalar potential generates the adjoint vevs in $\left\langle\varphi_{2}\right\rangle$ of eq (3.4). As an example for how this works, if $\Phi$ only appeared in powers of ${\rm Tr}\left(\Phi^{2}\right)$ in the numerator, then minimization of the scalar potential for the vacuum would try to make that trace vanish by imposing ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle=-{\rm tr}\left\langle\varphi_{1}^{2}\right\rangle$. Since ${\rm tr}\left\langle\varphi_{1}^{2}\right\rangle\neq 0$ (see the beginning of section 3), this tries to make ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle\neq 0$ for the vacuum solution.
Once the scalar vevs have been adjusted and a minimum of the quantum potential has been achieved, quantum superpotential terms can also produce fermion and scalar mass terms. For example, the term $O_{0}$ generates the following terms that modify the tree-level masses $m_{m11}$ and $m_{m22}$:
$-\tfrac{1}{3}\bar{O}_{0}\sum_{m}\left(\frac{\tilde{Q}_{m1}Q_{m1}}{\bar{\tilde{\phi}}_{21}\bar{\phi}_{21}}+\frac{\tilde{Q}_{m2}Q_{m2}}{\bar{\tilde{\phi}}_{22}\bar{\phi}_{22}}\right),$ (A.10)
where terms involving $1/\bar{\phi}_{23}$ and $1/\bar{\tilde{\phi}}_{23}$ have been ignored due to their inverse unification-scale vevs.
In addition, the $O_{0}$ part of the quantum potential also generates a Majorana mass term along with its scalar mass counterpart:
$\tfrac{4}{9}\bar{O}_{0}\left(\frac{\theta\tilde{\nu}_{W}^{\left(1\right)}}{\bar{\tilde{\phi}}_{21}}+\frac{\theta\nu_{W}^{\left(2\right)}}{\bar{\phi}_{22}}+\frac{\theta\tilde{\nu}_{W}^{\left(2\right)}}{\bar{\tilde{\phi}}_{22}}+\frac{\theta\nu_{W}^{\left(1\right)}}{\bar{\phi}_{21}}\right)^{2}$ (A.11)
$+$ $\tfrac{4}{9}\bar{O}_{0}\theta^{2}\left(\frac{(\tilde{f}_{21})_{1}}{\bar{\tilde{\phi}}_{21}}+\frac{(f_{22})_{2}}{\bar{\phi}_{22}}\right)\left(\frac{(\tilde{\phi}_{21})_{1}}{\bar{\tilde{\phi}}_{21}}+\frac{(\phi_{22})_{2}}{\bar{\phi}_{22}}\right)+...$
Since the vevs $\bar{\tilde{\phi}}_{21}$ and $\bar{\phi}_{22}$ are very small, the first line generates a very large Majorana mass $\tilde{M}_{\nu}$. The second line generates the corresponding scalar mass, where the notation $(\tilde{\phi}_{21})_{1}$ refers only to the first SU(3) component of that scalar.
Due to the relative magnitudes of the vevs, it is a good approximation to treat the above Majorana neutrino as being primarily for the combination $\tilde{\nu}_{M}=\cos\theta_{12}\tilde{\nu}_{W}^{\left(1\right)}+\sin\theta_{12}\nu_{W}^{\left(2\right)}$, where $\tan\theta_{12}=\bar{\tilde{\phi}}_{21}/\bar{\phi}_{22}$.
Another expression that is supergauge invariant is
${\rm Tr}\left(W^{0\alpha}W^{\prime}_{\alpha}\Phi\right)$ (A.12)
The above expression has dimension 4 and no axial anomaly charge, so it can replace $\left({\rm Tr}\left(\Phi^{2}\right)\right)^{2}$ in $O_{0}$. Since from eq (3.10), $\left\langle W^{0\alpha}\right\rangle\neq 0$, eq (A.12) generates quark mass terms that mix gaugino quarks with adjoint quarks, such as $\tilde{M}_{G}^{(1)}$ in eq (3.12).
The quantum superpotential can also include expressions like $O_{0}$ of eq (A.8) where the “meson” factors in the denominator are generalized in the following way:
$M^{\left(0\right)}_{m_{1}m_{2}FF^{\prime}}\to M^{\left(n\right)}_{m_{1}m_{2}FF^{\prime}}.$ (A.13)
These terms must have corresponding additional factors of $\Phi^{2}$ in the numerator. For example, a term where the flavor 2 and 3 mesons have a total of 3 additional factors of $\Phi^{2}$ in the denominator, there must be an additional factor of $\Phi^{2}$ in the numerator. One of these terms can have ${\rm Tr}\left(W^{0\alpha}W^{\prime}_{\alpha}\Phi^{3}\right)$ in the numerator. This term generates an adjoint-gaugino mass term for down-type quarks but not for up-type quarks (since the “11” component of $\langle\Phi^{2}\rangle$ is zero). The presence of terms like these provide justification for not forcing these types of terms to be the same in the up-type and down-type quark matrices of section 3. Terms involving ${\rm Tr}\left(\Phi^{4}\right)$ in the numerator can similarly lead to differences in adjoint-adjoint quark mass terms for up-type and down-type quarks.
Superpotential terms with ${\rm Tr}(\Phi^{2})^{3}$ in the numerator can lead to mass terms like that of eq (A.10), but where $\tilde{Q}_{m1}Q_{m1}\to\tilde{Q}_{m1}\langle\Phi^{2}\rangle Q_{m1}$. These terms only generate masses for the second two components of lepton triplets. This shows that the first component may have a different mass than the second two components. This freedom is used in the lepton mass matrices of section 3.
For superpotential terms in which the numerator has dimension 5 (such as $W^{0\alpha}W^{0}_{\alpha}{\rm Tr}\left(\Phi^{2}\right)$), there must be an odd number of $\Phi$ factors in the denominator. A term like this involving $M^{(1)}_{2111}$ generates a quark mass term for $\tilde{u}_{W}^{A}u^{(1)}$ that supplements the tree level term involving $\bar{\tilde{\phi}}_{21}\Gamma_{211}$. Similarly, many other quantum-generated couplings can supplement equivalent tree-level couplings.
Below the unification scale, it is possible to consider an effective theory where the $m=2$ SU(3) group has been broken down to SU(2). From section 4 of this paper, below the scale of $\tilde{M}_{\nu}$ (the heavy Majorana neutrino), the SU(2) group has a value of $b_{2}=-\tfrac{1}{6}$ for its running coupling. Following the logic earlier in this appendix, the following “meson” factors are gauge invariant to the SU(2), $U(1)_{Y}$ and $U(1)_{Y^{\prime}}$ groups and also satisfy the axial anomaly selection rule:
$M_{mm^{\prime}FF^{\prime}}^{\prime\left(n\right)}=\Lambda^{-1/18}_{\rm SU(2)}\tilde{Q}_{mF}T_{2}^{0}\Phi^{n}T_{2}^{0}Q_{m^{\prime}F^{\prime}}.$ (A.14)
Due to the factors of $T_{2}^{0}$ (from eq (1.7)), the above terms pick out leptons and scalars that only experience the SU(2) interaction, not the SU(3) strong interaction. The factor of $1/18$ is due to $b_{2}$ in the effective theory being $1/18$ the value of $b_{3}$ of eq (A.2) and the fact that all three flavors of leptons are still in the effective theory.
Since SU(2) is a self-conjugate group, there are additional terms in the SU(2) sector. Consider the following definitions for the $F=1,2$ fields in that sector:
$Q^{\prime}_{F}$ $=\left(\begin{array}[]{c}(\phi_{2F})_{1}+(\theta^{2}f_{2F})_{1}+\sqrt{2}\theta(\psi_{2F})_{1}\\ (\phi_{2F})_{2}+(\theta^{2}f_{2F})_{2}+\sqrt{2}\theta(\psi_{2F})_{2}\\ (\tilde{\phi}_{2F})_{3}+(\theta^{2}\tilde{f}_{2F})_{3}+\sqrt{2}\theta(\tilde{\psi}_{2F})_{3}\end{array}\right)$ (A.15)
$\tilde{Q}^{\prime T}_{F}$ $=\left(\begin{array}[]{c}(\tilde{\phi}_{2F})_{1}+(\theta^{2}\tilde{f}_{2F})_{1}+\sqrt{2}\theta(\tilde{\psi}_{2F})_{1}\\ (\tilde{\phi}_{2F})_{2}+(\theta^{2}\tilde{f}_{2F})_{2}+\sqrt{2}\theta(\tilde{\psi}_{2F})_{2}\\ (\phi_{2F})_{3}+(\theta^{2}f_{2F})_{3}+\sqrt{2}\theta(\psi_{2F})_{3}\end{array}\right),$
where the lower index outside of parentheses in each element refers to the original fundamental SU(3) index of that field. In $Q^{\prime}_{F}$ and $\tilde{Q}^{\prime}_{F}$, the third elements (the SU(2) singlets) have been changed to conjugate fields. The following terms are gauge invariant to the SU(2), $U(1)_{Y}$ and $U(1)_{Y}^{\prime}$ groups while also satisfying the axial anomaly selection rule:
$\Lambda^{-1/18}_{\rm SU(2)}\left(\tilde{Q}_{F}\varphi_{2}T^{\prime 2}\tilde{Q}^{\prime T}_{F}+\tilde{Q}^{\prime}_{F}T^{\prime 2}\varphi_{2}^{T}\tilde{Q}^{T}_{F}\right)$ (A.16)
$\Lambda^{-1/18}_{\rm SU(2)}\left(Q^{\prime T}_{F}T^{\prime 2}\varphi_{2}Q_{F}+Q^{T}_{F}\varphi_{2}^{T}T^{\prime 2}Q^{\prime}_{F}\right),$
where the 3$\times$3 matrix $T^{\prime 2}$ is like $t^{2}$ (half the Pauli matrix), except with $\tfrac{1}{2}$ in the “33” element. Also $Q_{F}$ (with no prime) in the above equation represents the bottom 3 components of the 6-vector $Q_{1F}+Q_{2F}$, and $\tilde{Q}_{F}$ is defined the same way.
Following similar logic to that of eqs (A.10) and (A.11), the superpotential terms of eq (A.16) will lead to flavor-1 charged lepton masses proportional to $c/\bar{\tilde{\phi}}_{21}^{2}$ and $1/\bar{\phi}_{21}^{2}$ for the Omega lepton and electron, respectively. In other words, it is proposed that the vevs generate the ratio
$m_{\Omega}:m_{e}=c/\bar{\tilde{\phi}}_{21}^{2}:1/\bar{\phi}_{21}^{2}\sim 10^{6}-10^{9}:1$ (A.17)
for mass terms $m_{\Omega}\tilde{e}_{W}^{(1)}e^{(1)}$ and $m_{e}\tilde{e}^{(1)}e_{W}^{(1)}$. The factor of $c$ is from eq (3.4). The range of multiples is to accommodate the wide range of possible Omega lepton masses described in section 3.
Due to the structure of eq (A.15), for each pair of two 2-component leptons that combine to form a Dirac mass, there are two scalar fields with that same mass contribution. In particular, the same mechanism that generates a mass of $m_{\Omega}$ for the Omega lepton generates that same mass contribution to the scalar fields $(\phi_{21})_{3}$ and $(\tilde{\phi}_{21})_{2}$. The total mass of these scalar fields can also include contributions from the scalar mass parameters $m_{21}$ and $\tilde{m}_{21}$ in the Lagrangian. Since $m_{21}$ should be significantly smaller than $m_{\Omega}$, the mass of $(\phi_{21})_{3}$ should be similar to that of the Omega lepton.
Presumably, detailed nonperturbative calculations could be performed to determine exact forms of superpotential terms for this theory. In the absence of those calculations which would produce the exact parameter values and scalar vevs at the quantum minimum, this paper has phenomenologically determined the values those parameters and vevs would have to take in order to reproduce experimental data.